### v1

# A linear-time algorithm computing the resident fitness in interacting trajectories

Abstract: The notion of a system of interacting trajectories was recently introduced in [13]. Such a system of $[0,1]$-valued piecewise linear trajectories arises as a scaling limit of the system of logarithmic subpopulation sizes in a certain population-genetic model (more precisely, a Moran model) with mutation and selection. By definition, the resident fitness is initially 0 and afterwards it increases by the ultimate slope of each trajectory that reaches height 1.
We show that although the interaction of $n$ trajectories may yield $\Omega(n^{2})$ slope changes in total, the resident fitness (at all times) can be computed algorithmically in $O(n)$ time. Our algorithm is given in terms of the so-called continued lines representation of the system of interacting trajectories. In the special case of Poissonian interacting trajectories where the birth times of the trajectories form a Poisson process and the initial slopes are random and i.i.d., we show that even the expected number of slope changes grows only linearly in time.
Keywords: (Poissonian) interacting trajectories, continued lines representation, algorithmic construction, resident fitness, speed of adaptation.
1 Introduction
The concept of a system of interacting trajectories has been recently introduced in [13]. Such a system consists of continuous and piecewise linear trajectories in a 2D coordinate system, where the $x$ axis represents time and the $y$ axis represents a magnitude that corresponds to the size of subpopulations of a large population of constant total size on a logarithmic scale. The value of the trajectories is in $[0,1]$ at each time. Each trajectory has a birth time and an initial slope. Until the birth time, the trajectory is constant zero, then it starts growing linearly with its initial slope. When a trajectory reaches height 1, it becomes constant 1, while the slope of all other trajectories currently at a positive height becomes reduced by the ultimate slope of this trajectory before reaching height 1. The slope of each trajectory is nonincreasing, and trajectories that reach height 0 again stay constant 0 forever. This interactive dynamics models the exponential growth/decay of mutant subpopulations in a certain stochastic population-genetic model, more precisely a Moran model, with mutation and selection, under a logarithmic scaling. Here, selection is strong and the mutation rate is in the so-called Gerrish–Lenski regime (dating back to [12]), see Remark 4 below.
A special case is the system of Poissonian interacting trajectories (PIT), where birth times are random, forming a Poisson process, and the positive initial slopes are random and i.i.d. (i.e. independent and identically distributed) and independent of the birth times. Under a suitable scaling of parameters, the family of logarithmic subpopulation sizes indeed converges to the PIT in distribution (in a suitable space with a suitable topology, see [13, Theorem 2.3]). In the PIT, the resident fitness $F(t)$ is defined as the sum of the ultimate slope of trajectories reaching height 1 up to time $t$. If the distribution of initial slopes has a finite first moment, $F(t)/t$ converges almost surely to a deterministic number $\overline{v}$ called the speed of adaptation, and it tends to infinity otherwise (cf. [13, Theorem 2.4]). Moreover, if the initial slopes have a finite second moment, then the resident fitness process satisfies a functional central limit theorem (see [13, Theorem 2.7]). The proofs of the latter two assertions are based on a renewal argument.
A closed formula for $\overline{v}$ is only known in the case when the initial slopes are deterministic and constant. In that case, the speed was already computed in [1]; this result was put into the context of the PIT by [13, Proposition 2.5]. On the other hand, one can clearly see the computation of the value of $F(t)/t$ for $t$ large as a Monte Carlo simulation for the value of $\overline{v}$. In the case when the fitness advantages have finite variance, we also know from the functional central limit theorem that typical fluctuations of $F(t)/t$ around $\overline{v}$ are of order $1/\sqrt{t}$. Thus, in order to approximate $\overline{v}$, it is useful to find an efficient algorithm for the simulation of the resident fitness of (Poissonian) interacting trajectories.
The main result of the present paper is that there is a deterministic algorithm that computes the resident fitness for all times in the case of a system of $n$ interacting trajectories (with deterministic and birth times and initial slopes, where the birth times have been sorted in advance) in $O(n)$ time. The existence of such a linear-time algorithm is a priori not obvious because as we will show in Proposition 9 below, the interactions between the trajectories may yield $\Omega(n^{2})$ slope changes in total. Our general conclusion is that computing the resident fitness at all times is in general easier than determining the slopes of all piecewise linear trajectories in all pieces. Our algorithm is described in terms of the so-called continued lines representation, a process introduced in [13] that is in one-to-one correspondence with the PIT but consists of rays (half-lines) instead of piecewise linear broken lines.
Moreover, our Proposition 12 below states that in the case of the Poissonian interacting trajectories, the expected number of slope changes up to time $t$ is $O(t)$ as $t\to\infty$. The proof of the latter assertion is based on a modified version of the aforementioned renewal argument of [13] combined with classical large-deviation estimates. This proof clearly shows that although the worst-case interaction between $n$ trajectories may provide $\Omega(n^{2})$ slope changes in total, scenarios where there are more than a sufficiently large constant times $t$ slope changes up to time $t$ are exponentially unlikely in the Poissonian case.
Let us note that apart from the PIT, there are many other recent examples of piecewise linear limiting processes arising in stochastic population-genetic [7] and population-dynamic [2, 3, 4, 5, 6, 10, 11, 15] models under a logarithmic scaling. These processes originate from models with higher mutation rates than the one of [13], under a different scaling of time. The limiting processes are typically deterministic and their dynamics is also a bit different from the one of the interacting trajectories, thanks to multiple mutations between two given types of individuals and also due to other biological phenomena depending on the model (e.g. horizontal gene transfer, asymmetric competition, dormancy etc.). It is possible that variants or extensions of our algorithm can be used for simulating these limiting processes efficiently, but due to length constraints, we defer such questions to future work.
The rest of this paper is organized in follows. In Section 2 we recall the definition of a system of (Poissonian) interacting trajectories from [13] and we provide a short summary of its biological motivation. In Section 3 we present our results and in Section 4 we prove them. Finally, in Section 5 we briefly discuss some consequences of our assertions and proofs.
2 Setting
The definition of a system of (Poissonian) interacting trajectories that we will use in the present paper is a somewhat simplified version of the definition appearing in [13, Sections 2.1 and 3.1] (see Remark 7 below for the differences between the two definitions). It is given as follows.
Writing =$\mathbb{N}=\{1,2,\ldots\}$ and $\mathbb{N}_{0}=\{0,1,2,\ldots,\}$, for $n\in\mathbb{N}_{0}$ we define $[n]=\{1,2,\ldots,n\}$, moreover we put $[\infty]=\mathbb{N}$. We denote the space of continuous and piecewise linear trajectories $h$ from $[0,\infty)$ to $[0,1]$ such that $h(t)\in[0,1]$ for all $t\geq 0$ by $\mathcal{C}_{\rm PL}$. Each $h\in\mathcal{C}_{\rm PL}$ has at time $t$ a height $h(t)$ and a (right) slope
$v_{h}(t):=\lim\limits_{\delta\downarrow 0}\tfrac{1}{\delta}(h(t+\delta)-h(t)).$ (1)
Assume that for some $n\in\mathbb{N}_{0}\cap\{\infty\}$ we are given a configuration of pairs $(t_{i},a_{i})$ consisting of a birth time $t_{i}$ and an initial slope $a_{i}$
$\beth=((t_{i},a_{i}))_{1\leq i<n+1}\in([0,\infty)\times(0,\infty))^{[n]},$ (2)
with $0\leq t_{1}<t_{2}\cdots<t_{n}$ if $n<\infty$ and with $0\leq t_{1}<t_{2}<\ldots$ and $\lim_{n\to\infty}t_{n}=\infty$ if $n=\infty$ (where we use the convention $\infty+1=\infty$). $\beth$ (beth) specifies that the trajectory $h_{i}$, $11\leq i<n+$ 1, has height 00 for $t\in[0,t_{i}]$ and right slope $a_{i}$ at its birth time $t_{i}$.
Definition 1 (System of interacting trajectories, [13])
For $\beth$ as in (2), let
$\mathbb{H}=(h_{i})_{0\leq i<n+1}\in\big{(}\mathcal{C}_{\rm PL}\big{)}^{[n]\cup\{0\}}$ (3)
result from the following deterministic interactive dynamics on $(\mathcal{C}_{\rm PL})^{[n]\cup\{0\}}$ :
  • One initial trajectory $h_{0}\in\mathcal{C}_{\rm PL}$ starts at height $h_{0}(0)=1$ with slope $v_{h_{0}}(0)=0$. All other trajectories $h_{i}$, $11\leq i<n+$ 1 start at height $h_{i}(0)=0$ with slope $v_{h_{i}}(0)=0$.
  • Trajectories continue with constant slope until the next birth time is reached or one of the trajectories reaches either 1 from below or 0 from above.
  • For $t\geq 1$, at the birth time $t_{i}$ the slope of trajectory $h_{i}$ jumps from 00 to $a_{i}$.
  • Whenever at some time $t$ at least one trajectory reaches height $1$ from below, the slopes of all trajectories whose height is in $(0,1]$ at time $t$ are simultaneously reduced by
    $v^{\ast}:=\max\limits\{v_{h}(t-)\mid h\in\mathbb{H}\mbox{ such that }h(t)=1\},$
    i.e. for all $h\in\mathbb{H}$ with $h(t)>0$
    $v_{h}(t):=v_{h}(t-)-v^{\ast}.$ (4)
  • Whenever a trajectory at some time $t$ reaches height 00 from above, its slope is instantly set to 00, and this trajectory then stays at height 00 forever.
    Let $e_{i}$ denote the time when this happens to the $i$-th trajectory (where we set $e_{i}=\infty$ if it never happens) where $10\leq i<n+$ 1. For $t\geq 0$ we say that the $i$-th trajectory is alive at time $t$ if $t_{i}\leq t<e_{i}$.
We call this $\mathbb{H}$ the system of interacting trajectories initiated by $\beth$, and denote it by $\mathbb{H}(\beth)$. An alternative, equivalent representation of the system of interacting trajectories is the so-called continued lines representation, which plays an important role in the proof of Theorem 11 and will be introduced in Remark 8 below. Instead of broken lines, this representation consists of nondecreasing half-lines starting at the birth times of the corresponding trajectories, which makes it more amenable for algorithmic investigations.
Definition 2 (Resident change times, resident type, and resident fitness, [13])
Let $\mathbb{H}$ be as in (3), following the dynamics specified in Definition 1.
  • The times at which one of the trajectories $h_{i}$, $i>0$, reaches height $1$ from below will be called the resident change times.
  • For $t>0$ we call
    $\varrho(t)=\varrho^{(\beth)}(t):=\arg\max\{v_{h_{i}}(t-)\mid 0\leq i<n+1,h_{i}(t)=1\},$ (5)
    the resident type at time $t$.
  • With the definition $f(0)=0$, we define the resident fitness $f(t)$, $t\geq 0$, by decreeing that $f$ at any resident change time $r$ has an upward jump with
    $f(r)-f(r-)=\max\{v_{h}(r-)\mid h\in\mathbb{H},\,h(r)=1\}$ (6)
and remains constant between any two subsequent resident change times.
The following remark is an adaptation of [13, Remark 3.4] to our situation.
Remark 3 ([13])
  1. 1.
    The resident fitness $f$ obeys
    $f(t)=a_{\varrho(t)}+f(t_{\varrho(t)}),\quad t\geq 0.$ (7)
  2. 2.
    If $t\geq 0$ is not a resident change time, then $\varrho(t)$ is the unique $i>0$ for which $h_{i}(t)=1$. If $t$ is a resident change time and $h_{i}$ is the only trajectory that reaches height 1 from below at time $t$, then $\varrho(t)=i$. More generally, $\varrho(t)$ in (5) is the fittest of all types that are at height $1$ at time $t$.
  3. 3.
    The resident type $\varrho(t)$ defined in (5) satisfies (with $t_{0}:=0$)
    $\varrho(t)=\max\{i\colon 0\leq i<n+1,\,h_{i}(t)=1\}$ (8)
    (see [13, Remark 3.4] for an explanation).
Remark 4
The biological interpretation of this interactive dynamics is as follows (see [13] for more details). We consider a Moran model with mutation and selection. This population-genetic model is a continuous-time Markov chain. The total population size $N$ of the model is fixed, and the interacting trajectories arise as a scaling limit as $N\to\infty$. The height $h_{i}(t)$ of a given trajectory at a given time $t$ corresponds to the logarithmic size of a subpopulation of genetically identical individuals divided by $\log N$ in the Moran model. Thus, the height of a non-extinct subpopulation is always between 0 and 1, and upon extinction, we cut the trajectory at height 0 by definition, so that it cannot become negative.
At time 0, all $N$ individuals belong to the initial resident population and have the same fitness value 0. Beneficial mutations occur in a randomly chosen individual at time $t_{i}\log N$, with a positive fitness advantage $a_{i}$, for all $11\leq i<n+$ 1. Mutants have a random selective advantage of strength not depending on $N$ (this is why selection is called strong), allowing them to reproduce faster. Mutant subpopulations tend to either go extinct rapidly or grow approximately exponentially, hence their logarithmic population size divided by $\log N$ grows approximately linearly and takes $\Theta(\log N)$ time to reach the vicinity of 1.
The per generation mutation rate is $\mu_{N}=\Theta(1/\log N)$, in other words, the model belongs to the so-called Gerrish–Lenski mutation regime. A characteristic property of this regime is that the waiting time between two consecutive mutations is of the same order of magnitude (in our case it is $\Theta(\log N)$) as the time until a mutant subpopulation reaches height 1. Hence, typically multiple mutant families are fighting with each other for survival, which is a phenomenon referred to as clonal interference.
Speeding up time by a factor of $\log N$, we obtain approximately the system $\mathbb{H}(\beth)$; the corresponding convergence result is [13, Theorem 2.3]. Each mutation happens in an individual of the current resident population with high probability. Hence, the limiting trajectory $h_{i}$ starts with slope $a_{i}$ at time $t_{i}$ and keeps this slope until some subpopulation becomes resident (i.e. its trajectory reaches height 1). Upon this resident change, the slope of each subpopulation currently at a positive height drops by the increment of the resident fitness, i.e. the ultimate slope of the trajectory of the new resident before reaching height 1 (cf. (4)). Thus, slopes may become negative, corresponding to an approximately exponential decay of subpopulations that are already unfit compared to the current resident. The time when a trajectory hits zero marks the extinction time of the corresponding mutant subpopulation in the Moran model.
The authors of [13] were mainly interested in the case of random $\beth$ where $n=\infty$ and
  • the role of $(t_{i})_{i\in\mathbb{N}}$ is played by the sequence of arrival times $(T_{i})_{i\in\mathbb{N}}$ of a Poisson process $(T_{i})_{i\in\mathbb{N}}$ with intensity $\lambda>0$ (i.e. putting $T_{0}=0$, $T_{i}-T_{i-1}$, $i=1,2,\ldots$ are i.i.d. exponentially distributed random variables with parameter $\lambda$),
  • whereas the role of $(a_{i})_{i\in\mathbb{N}}$ is played by a sequence $(A_{i})_{i\in\mathbb{N}}$ of i.i.d. strictly positive random variables that is independent of the Poisson process $(T_{i})_{i\in\mathbb{N}}$, where the common marginal distribution of each $A_{i}$ on $(0,\infty)$ is denoted by $\gamma$ (i.e. $\gamma$ is a Borel probability measure on $(0,\infty)$ such that for any $x\in\mathbb{R}$, $\gamma((-\infty,x])=\mathbb{P}(A_{1}\leq x)$).
In this case, we call $\mathbb{H}(\beth)$ the system of Poissonian interacting trajectories (or PIT for short) with parameters $(\lambda,\gamma)$, or briefly as the $\mathrm{PIT}(\lambda,\gamma)$. The corresponding trajectories will also be denoted by capital letters $H_{i}$ (note that they are deterministic functions of the random input $((T_{j},A_{j}))_{j\in\mathbb{N}}$). The resident fitness $F$ that we mentioned in the introduction corresponds to $f$ from Definition 5 in the case of the PIT. We will generally exclude conditions that are degenerate according to the following definition.
Definition 5
Let $\beth$ be as in (2). We say that $\beth$ is degenerate if the corresponding $(h_{i})_{i\in[n]\cup\{0\}}$ defined according to Definition 1 satisfies at least one of the following conditions:
  • There exists $t>0$ and three pairwise distinct indices $i,j,k\in[n]\cap\{0\}$ such that $h_{i}(t)=h_{j}(t)=h_{k}(t)=1$.
  • There exists $j\in[n]\cap\{0\}$ such that the extinction time of $h_{j}$, i.e. $\inf\{t>t_{0}\mid h_{j}(t)=0\}$, is a resident change time or equals $t_{i}$ for some $i\in[n]$.
  • There exists $i\in[n]$ such that $t_{i}$ is a resident change time.
Elementary properties of the Poisson process imply the following claim.
Claim 6
[…]
Remark 7
Apart from considering only finitely many trajectories instead of countably many, the main changes in the above definitions compared to their analogues in [13] are the following:
  • We ignore general initial conditions with multiple trajectories at positive heights at time 0, which [13] used in proofs. For simplicity we allow $t_{1}=0$ (and not only $t_{1}>0$).
  • The original definition of the PIT includes trajectories whose slope and height stays zero forever. These describe mutant families going extinct rapidly by chance shortly after the mutation. This possibility is biologically relevant but algorithmically uninteresting.
Remark 8 (The continued lines representation)
An alternative, equivalent representation of the PIT, called the continued lines representation, was provided in [13, first arXiv version, Section 5.3]. [^1] This section is only available in the first and second version of [13] on arXiv. This representation can be simulated without computing the resident fitness at any time and without any calculations involving already decaying height functions, which is useful for algorithmic investigations. We describe this representation for a general system of interacting trajectories $\mathbb{H}(\beth)$ as follows.
At any time $t\geq 0$ we consider a finite family of rays (half-lines) $\{l_{0},l_{1},\ldots,l_{k}\}$ that are infinite towards the right, where $l_{0}$ is the constant 1 ray starting at time 0, and we denote by $l^{r}(t)=\max\{l_{1}(t),\ldots,l_{k}(t)\}$ the pointwise maximum, namely the resident ray, with $f(t)$ defined in (2) being equal to its slope at time $t$. When the initial configuration $\beth$ delivers a new point $(t_{k+1},a_{k+1})$, we add a line $l_{k+1}$ to the system starting at $l^{r}(t_{k+1})-1$ with slope $f(t_{k+1})+a_{k+1}$. Notably, this gives an alternative construction of $\mathscr{H}$ as for any $k$, $h_{k}(t)=l_{k}(t)-(l^{r}(t)-1)$. After $t_{k}$, the $k$-th trajectory of $\mathbb{H}(\beth)$ hits 0 when $l_{k}(t)$ reaches $l^{r}(t)-1$ again.
Let us assume that $\mathbb{H}(\beth)$ is nondegenerate. Then, for $k\geq 0$, if $j\in\{0,1,\ldots,k\}$ satisfies $l^{r}(t_{k+1})=l_{j}(t_{k+1})$, we say that $j$ is the parent of $k+1$ (note that there is a unique such $j$ is well-defined under the nondegeneracy assumption).
3 Results
For a sequence $(\beth^{(n)})_{n\in\mathbb{N}}$ of non-degenerate initial data $\beth^{(n)}=((t_{i}^{(n)},a_{i}^{(n)}))_{1\leq i<n+1}$, writing $\mathbb{H}(\beth^{(n)})=(h_{i}^{(n)})_{0\leq i<n+1}$, we denote by
$k_{n}(\beth^{(n)})$ $=\sum_{t\geq 0\colon t\text{ resident change time of }\mathbb{H}(\beth^{(n)})}\sum_{i=0}^{n}\mathds{1}_{\{h_{i}^{\prime(n)}(t-)\neq h_{i}^{\prime(n)}(t+)\}\cap\{h_{i}^{(n)}(t)>0\}}$
$=\sum_{t\geq 0\colon t\text{ resident change time of }\mathbb{H}(\beth^{(n)})}\big{|}\{i\in\{0,1,\ldots,n\}\colon h_{i}\text{ is alive at time }t\}\big{|}$
the total number of kinks (slope changes) in the system of interacting trajectories $\mathbb{H}(\beth^{(n)})$. Then we have the following result.
Proposition 9
There exists a sequence $(\beth^{(n)})_{n\in\mathbb{N}}$ as above such that $k_{n}(\beth^{(n)})=\Omega(n^{2})$.
The proof of Proposition 9 can be found in Section 4.1.
Remark 10
Apart from the slope changes corresponding to the definition of $k_{n}$, at each time $t_{i}$, $i=1,\ldots,n$, there is an additional slope change where the slope of the $i$-th trajectory switches from 00 to $a_{i}$. Moreover, at most $n$ trajectories eventually hit 0 again and then their slope drops to 0. Altogether, these are at most $2n$ additional slope changes, and precisely $2n$ unless there are multiple indices $j$ with the same value of $f(t_{j})+a_{j}$ as the ultimate resident. (In the case of the PIT, the latter is only possible if $\mathbb{P}(A_{i}=c)>0$ holds for some $c>0$.) Hence, given Proposition 9, ignoring these $O(n)$ additional slope changes will not yield substantial changes in the worst-case estimates of the number of kinks.
Given Proposition 9, computing the slopes of all trajectories in a system of $n$ interacting trajectories clearly takes $\Omega(n^{2})$ time. One might think that even computing the resident fitness for all times takes a quadratic amount of time. However, we have the following theorem, where $\beth$ counts as ordered if the corresponding $t_{i}$’s are sorted in the input and unordered otherwise.
Theorem 11
The algorithmic complexity of finding $t\mapsto(\varrho^{(\beth)}(t))_{t\geq 0}$ for any finite, nondegenerate, ordered $\beth$ of size $n$ is $\Theta(n)$, and it is $\Theta(n\log n)$ in the unordered case.
An algorithm proving the theorem will be provided in Section 4.2. The key idea is to use the continued lines representation, which makes it possible to provide such an algorithm using a simple data structure (involving two pointers and a doubly linked list), despite the possibly large number of kinks.
The output of the algorithm is a list containing the time intervals of residencies in consecutive order, together with the indices of the residents and the corresponding resident fitness values. Given this output, in Section 5.1 we explain how to determine the set of trajectories that reach fixation and the ancestral relations between these trajectories (the notion of fixation will also be explained there), whereas in Section 5.2 below we summarize how to compute the resident fitness at a given time and the trajectory $(h_{i}(t))_{t\geq 0}$ for fixed $i$ efficiently.
Our next proposition shows that if one considers the PIT where the birth times of trajectories form a Poisson process, then even the expected number of kinks will only grow linearly in time. (Due to the time-continuous nature of the Poisson process, here one has to work with a notion of kinks (3) slightly different from the one (3) above; see Remark 13 for a related discussion.)
Let $\lambda,\gamma$ as above and let us consider the $\mathrm{PIT}(\lambda,\gamma)$. For $t\geq 0$ we write $N_{t}=\{i\geq 1\colon T_{i}\leq t\}$ where $T_{i}$ is the birth time of the $i$-th trajectory (with the aforementioned convention that $T_{0}=0$), then $N_{t}$ is $\mathrm{Poisson}(\lambda t)$ distributed for any $t>0$, and for any $n\in\mathbb{N}$ and $0=t_{0}<t_{1}<\ldots<t_{n}$, $N_{t_{i}}-N_{t_{i-1}}$, $i=1,\ldots,n$ are independent). Define the total number of kinks up to $t$ as
$K_{t}=\sum_{0<s\leq t\colon s\text{ resident change time}}\ \sum_{i\in\mathbb{N}_{0}}\mathds{1}_{\{H_{i}^{\prime}(t-)\neq H_{i}^{\prime}(t+)\}}.$
Proposition 12
We have $\mathbb{E}(K_{t})=O(t)$.
The proof of Proposition 12 will be carried out in Section 4.3. This proof is based on an extension of the aforementioned renewal argument of [13], combined with classical large-deviation inequalities. The main message of this proof is that the clustering of trajectories that leads to many kinks (like in the construction in the proof of Proposition 9) is exponentially unlikely in the number of participating trajectories if these trajectories follow each other directly in the PIT according to their birth times.
Remark 13 (Relation between Propositions 9 and 12)
$K_{t}$ is the number of resident changes up to time $t$ counted with multiplicity: Each resident change is counted as many times as many trajectories change their slopes due to this resident change. Note that putting $\beth^{(n)}=((T_{i},A_{i}))_{i\in[n]}$, the total number of kinks due to the interactions of the first $N_{t}$ trajectories is in general an upper bound for $K_{t}$, i.e.
$K_{t}\leq k_{N_{t}}(\beth^{(N_{t})}),$
the difference $k_{N_{t}}(\beth^{(N_{t})})-K_{t}$ being the number of those kinks in $\mathbb{H}(\beth^{(N_{t})})$ that happen after time $t$. However, it will be apparent from the proof of Proposition 12 that following time 0, after a stopping time with finite expectation, a state will be reached where $H_{i}(t)=0$ for all but one $i$. Hence, by the strong Markov property, this state will be visited infinitely often, with i.i.d. waiting times with finite expectation inbetween. Since a trajectory can only have kinks before its height ultimately reaches 00, it follows that the difference between $K_{t}$ and $k_{N_{t}}(\beth^{(N_{t})})$ is of finite order. This way, Proposition 12 can indeed be seen as complementary to Proposition 9 in the Poissonian case.
4 Proofs
4.1 Lower bound construction for the number of kinks: proof of Proposition 9
Proof: For $n\in\mathbb{N}$ describe an input $\beth^{(n)}$ so that $(\mathbb{H}(\beth^{(n)})_{n\in\mathbb{N}}$ satisfies the conditions of the proposition and we verify that it indeed satisfies them. For simplicity, we use the continued lines representation introduced in Remark 8 (which is in bijection with the representation as a system of interacting trajectories).
For $n\in\mathbb{N}$ and $k\in\{0,1,\ldots,n\}$, we define the $k$-th ray $l_{k}$ corresponding to the continued lines representation of $\mathbb{H}(\beth^{(n)})$ via the equations $l_{0}(t)=1$, $t\geq 0$, and for $k\geq 1$,
$l_{k}(t)=k\Big{(}t-(1-\frac{1}{k}+\frac{k-1}{n^{2}})\Big{)},\qquad t\geq 1-\frac{1}{k}+\frac{k-1}{n^{2}}.$
First we claim that $l_{k}$ and $l_{k+1}$ intersect at time $1+\frac{2k}{n^{2}}$. Indeed, this is clear for $l_{0}(t)=1$ and $l_{1}(t)=t$. Moreover, for $k\geq 1$, $t$ is the time when $l_{k}$ and $l_{k+1}$ intersect if and only if
$k\big{(}t-1+\frac{1}{k}-\frac{k-1}{n^{2}}\big{)}=(k+1)\big{(}t-1+\frac{1}{k+1}-\frac{k}{n^{2}}\big{)},$
which is satisfied if and only if $t=1+\frac{2k}{n^{2}}$. Since $l_{k+1}$ is steeper than $l_{k}$, this implies that
$l_{k+1}(t)>l_{k}(t)\quad\Leftrightarrow\quad t>1+\frac{2k}{n^{2}}.$
We now claim that it is even true that
$l_{0}\text{ is resident on }[0,1)\text{ and for $k\geq 1$, }l_{k}\text{ is resident on }\big{[}1+\frac{2(k-1)}{n^{2}},1+\frac{2k}{n^{2}}\big{]}.$ (10)
Indeed, recall that at a given time $t$, the index of the resident is the index of the ray that is at the highest position among all rays at time $t$ (which implies that the ray has already been started by time $t$, and in the case of a draw, the index with the higher slope is the index of the resident). Now, (10) follows inductively from the fact that 00 is obviously resident at time 00, type $k$ cannot be resident at time $t<1+\frac{2(k-1)}{n^{2}}$ because then $l_{k-1}(t)>l_{k}(t)$, and it cannot be resident at time $t>1+\frac{2k}{n^{2}}$ either because then $l_{k+1}(t)>l_{k}(t)$, but at any time there must be a resident type, so $l_{k}$ must be resident for $\frac{2(k-1)}{n^{2}}<t<\frac{2k}{n^{2}}$ and thus by construction also for $t=\frac{2(k-1)}{n^{2}}$.
Now we claim that the 00-th trajectory of $\mathbb{H}(\beth^{(n)})$ is still alive at time $1+\frac{2(n-1)}{n^{2}}$, equivalently, $l_{0}(1+\frac{2(n-1)}{n^{2}})>l^{r}(1+\frac{2(n-1)}{n^{2}})-1=l_{n}(1+\frac{2(n-1)}{n^{2}})-1$. Indeed, we have that
$l_{0}(t)=1,\qquad l_{n}(1+\frac{2(n-1)}{n^{2}})=n\big{(}1+\frac{2(n-1)}{n^{2}}-1+\frac{1}{n}-\frac{n-1}{n^{2}}\big{)}=n\big{(}\frac{2n-1}{n^{2}}\big{)}=2-\frac{1}{n}.$
By construction, for $t>1+\frac{2(n-1)}{n^{2}}$, $l_{0}(t)$ is definitely the smallest one among all values $l_{i}(t)$, $i\in\{0,1,\ldots,n\}$. This implies that all other trajectories of $\mathbb{H}(\beth^{(n)})$ are also alive at time $1+\frac{2(n-1)}{n^{2}}$. Since they are all already born (i.e. the corresponding rays of the continued lines representation are already started) before time $1$, it follows that at all resident change times all trajectories of $\mathbb{H}(\beth^{(n)})$ are alive and thus they change their slopes. This yields $k_{n}(\beth^{(n)})=n(n+1)=\Omega(n^{2})$.  $\Box$
4.2 Proof of Theorem 11
The algorithm for the ordered case processes births one at a time, in ascending order, while maintaining a list of residents. For trajectory $i$, the parent is found by linear search from the start of the resident list forwards, and whether it becomes resident is determined by linear search from the end of the resident list backwards.
Recall from Remark 8 that in the continued lines representation, each trajectory birth is a pair $(t_{i},a_{i})$, where $t_{i}$ is the birth time and $a_{i}$ is the slope relative to its parent.
A trajectory is a tuple $(t_{i},p_{i},b_{i},y_{i})$, where $t_{i}$ is the birth time, $p_{i}$ is the index of the trajectory of the parent, who is the resident at time $t_{i}$, $b_{i}=b_{p_{i}}+a_{i}$ is the slope, and $y_{i}$ is the initial height, which is one less than the height of the parent at time $t_{i}$. This corresponds to a ray starting at point $(t_{i},y_{i})$ with slope $b_{i}$. A residency is a pair $(s_{j},m_{j})$, where $s_{j}$ is the residency start time and $m_{j}$ is the trajectory index. A residency starts when its trajectory intersects the trajectory of the current resident, so the residencies form a convex broken line. The $j$th residency starts at point $(s_{j},y_{m_{j}}+b_{m_{j}}(s_{j}-t_{m_{j}}))$.
Initially, there is one trajectory in the system, $(t_{0}=0,p_{0}=-1,b_{0}=0,y_{0}=1)$, which is the current resident, $(s_{0}=0,m_{0}=0)$. In step $i$, the forward linear search finds the parent of trajectory $i$, which is the last resident $j$ with $s_{j}\leq t_{i}$. The $i$th trajectory is $(t_{i},p_{i}=m_{j},b_{i}=(b_{p_{i}}+a_{i}),y_{i}=(y_{p_{i}}+b_{p_{i}}(t_{i}-t_{p_{i}})-1))$. Since the input is ordered by $t_{i}$, the search can resume at the position it stopped at in the previous step.
If ray $i$ is below the convex broken line of residents, it does not become a resident. Otherwise, there is an intersection point, which is found by backward linear search. Starting from the last resident in the list, it discards all residency segments from the previous solution below the new ray until the intersection point $(x,y)$ is found and in that case the new residency segment $(x,i)$ is added to the end.
The list of residents may be stored in a doubly linked list. Since both linear searches pass each trajectory at most once, the algorithm has time $O(n)$. It is easy to see, that it is also $\Theta(n)$, for example when all trajectories become residents.
When the input is not ordered, it can be shown that the optimal time for solving the problem is $\Omega(n\log n)$, since for a given unordered list of pairwise distinct integers $\{x_{1},\dots,x_{n}\}$, the input trajectories $\{(x_{1},2),\dots,(x_{n},2)\}$ will result in a resident list containing all $x_{i}$, in sorted order. It is also $\Theta(n\log n)$ because the same algorithm with initial sorting can be used.
4.3 Proof of Proposition 12
Proof: Let $E_{0}=0<E_{1}<E_{2}<\ldots$ be the renewal extinction times, defined via
$E_{i}=\inf\{t\geq E_{i-1}\mid\text{ only the resident type at time $t$ has a positive height}\}$
for $i\geq 1$. Note that for $i\geq 1$, on the time interval $(E_{i-1},E_{i}]$ only trajectories born in $(E_{i-1},E_{i}]$ can become resident, and trajectories whose height function can suffer any kinks on this time interval due to resident changes are also the same trajectories plus the trajectory that was resident at time $E_{i-1}$. Indeed, any other trajectory born on $[0,E_{i-1}]$ goes extinct by the time $E_{i-1}$. Hence, for $t\geq 0$, putting $\eta_{t}=|\{i\geq 1\colon E_{i}\leq t\}|=\sup\{i\geq 0\colon E_{i}\leq t\}$, we have $E_{\eta_{t}}\leq t<E_{\eta_{t}+1}$. Note also that $\eta_{t}\leq N_{t}$ for all $t$ because at each $E_{i}$ a trajectory born before $E_{i}$ becomes resident (and one trajectory can only become resident at most once).
It can be proven analogously to the proof of [13, Lemma 5.5] that $E_{i}-E_{i-1}$ is stochastically dominated by a geometric random variable, and the random variables $(E_{i}-E_{i-1})_{i\geq 1}$ are i.i.d. (in other words, $(E_{n})_{n\in\mathbb{N}}$ is a renewal process). They have a finite first moment $m_{1}:=\mathbb{E}(E_{1})$ and a finite second moment $m_{2}:=\mathbb{E}(E_{1}^{2})$. This together with the tower property implies that
$\mathbb{E}(K_{E_{i}}-K_{E_{i-1}})$ $\leq\mathbb{E}((N_{E_{i}}-N_{E_{i-1}})(N_{E_{i}}-N_{E_{i-1}}+1))$
$=\mathbb{E}(\mathbb{E}((N_{E_{i}}-N_{E_{i-1}})(N_{E_{i}}-N_{E_{i-1}}+1)|E_{i}-E_{i-1}))=\mathbb{E}(\lambda^{2}(E_{i}-E_{i-1})^{2}+2\lambda(E_{i}-E_{i-1}))$
$=\lambda^{2}m_{2}+2\lambda m_{1}.$
Thus, for $n\geq 1$,
$\mathbb{E}(K_{E_{n}})=\sum_{i=1}^{n}\mathbb{E}(K_{E_{i}}-K_{E_{i-1}})\leq n(\lambda^{2}m_{2}+2\lambda m_{1}).$
The same arguments also imply the (much less tight) bound
$K_{t}\leq N_{t}(N_{t}+1)$
almost surely for all $t\geq 0$. Since $(E_{n})_{n\in\mathbb{N}}$ is a renewal process with increments having finite mean $m_{1}$, we have
$\lim_{n\to\infty}\frac{E_{n}}{n}=m_{1}$
almost surely thanks to the strong law of large numbers, and
$\lim_{t\to\infty}\frac{\eta_{t}}{t}=\frac{1}{m_{1}}$
almost surely (cf. [9, Theorem 2.5.10] for the latter assertion). Using Hölder’s inequality and (4.3), we conclude that for $\delta>0$, for all $t$ sufficiently large
$\mathbb{E}(K_{t})$ $\leq\mathbb{E}(K_{E_{\eta_{t}+1}}\mathds{1}_{\{\eta_{t}+1\leq(\frac{1}{m_{1}}+\delta)t\}})+\mathbb{E}(K_{t}\mathds{1}_{\{\eta_{t}+1>(\frac{1}{m_{1}}+\delta)t\}})$
$\leq\mathbb{E}(K_{E_{\lceil(\frac{1}{m_{1}}+\delta)t\rceil}})+\sqrt{\mathbb{E}(K_{t}^{2})}\sqrt{\mathbb{E}\Big{(}(\mathds{1}_{\{\eta_{t}+1>(\frac{1}{m_{1}}+\delta)t\}})^{2}\Big{)}}$
$\leq(\frac{1}{m_{1}}+\delta+1)t(\lambda^{2}m_{2}+2\lambda m_{1})+\sqrt{\mathbb{E}(K_{t}^{2})}\sqrt{\mathbb{P}\Big{(}\eta_{t}+1>(\frac{1}{m_{1}}+\delta)t\Big{)}}$
$\leq(\frac{1}{m_{1}}+\delta+1)t(\lambda^{2}m_{2}+2\lambda m_{1})+\sqrt{\underbrace{\mathbb{E}(N_{t}^{4}+2N_{t}^{3}+N_{t}^{2})}_{\in O(t^{4})}}\sqrt{\mathbb{P}\Big{(}\eta_{t}>(\frac{1}{m_{1}}+\delta)t-1\Big{)}}$
$\leq(\frac{1}{m_{1}}+\delta+1)t(\lambda^{2}m_{2}+2\lambda m_{1})+O(t^{2})\sqrt{\mathbb{P}\Big{(}\eta_{t}>(\frac{1}{m_{1}}+\delta)t-1\Big{)}}.$
Here, the assertion that $\mathbb{E}(N_{t}^{4}+2N_{t}^{3}+N_{t}^{2})\in O(t^{4})$ follows from the well-known fact that a Poisson distributed random variable $X$ with parameter $\lambda$ satisfies $\mathbb{E}(X(X-1)\ldots(X-k+1))=\lambda^{k}$ for any $k\geq 1$, from which one can easily derive that $\mathbb{E}(X^{4})=(1+o(1))\lambda^{4}$ as $\lambda\to\infty$.
To treat the last term on the right-hand side of (4.3), we use standard large-deviation estimates. First note that there exists $\delta^{\prime}>0$ not depending on $t$ such that
$\Big{\{}\eta_{t}>\big{(}\frac{1}{m_{1}}+\delta\big{)}t-1\Big{\}}=\Big{\{}E_{\lceil(\frac{1}{m_{1}}+\delta)t\rceil-1}\leq t\Big{\}}\subseteq\Big{\{}E_{\lceil(\frac{1}{m_{1}}+\delta)t\rceil-1}\leq(1-\delta^{\prime})m_{1}\big{(}\lceil(\frac{1}{m_{1}}+\delta)t\rceil-1\big{)}\Big{\}}.$
Since $\Lambda(\alpha):=\log\mathbb{E}(\mathrm{e}^{\alpha E_{1}})<\infty$ for $\alpha>0$ small (which also follows from the stochastic domination by a geometric random variable), Cramér’s theorem [8, Section 2.2] implies that if $\delta^{\prime}$ is sufficiently small, then
$\limsup_{n\to\infty}\frac{1}{n}\log\mathbb{P}\Big{(}E_{n}\leq(1-\delta^{\prime})m_{1}n\Big{)}\leq-\inf_{t\leq(1-\delta^{\prime})m_{1}}\Lambda^{*}(t)=:-I_{0}<0,$
where the rate function $\Lambda^{*}\colon\mathbb{R}\to[0,\infty]$ is given as
$\Lambda^{*}(a)=\sup_{\alpha\in\mathbb{R}}(\alpha a-\Lambda(\alpha)),$
and together with the fact that $\Lambda^{\prime}(0)=m_{1}$ and $\Lambda^{*}(m_{1})=0$, the strict negativity in (4.3) follows from the assertion of [8, Exercise 2.2.24] for $\delta^{\prime}>0$ small enough. We conclude that for all sufficiently large $t$,
$\sqrt{\mathbb{P}\Big{(}\eta_{t}>(\frac{1}{m_{1}}+\delta)t\Big{)}}\leq\mathrm{e}^{-\frac{1}{2}I_{0}(\lceil(\frac{1}{m_{1}}+\delta)t\rceil-1)}\leq\mathrm{e}^{-\frac{1}{2}I_{0}(\frac{1}{m_{1}}+\delta/2)t}.$
Thus, the right-hand side of (4.3) is bounded from above by
$(\frac{1}{m_{1}}+\delta+1)t(\lambda^{2}m_{2}+2\lambda m_{1})+O(t^{2})\mathrm{e}^{-\frac{1}{2}I_{0}(\frac{1}{m_{1}}+\delta/2)t}\in O(t),$
as wanted.  $\Box$
5 Discussion
5.1 Ancestral relations and fixations
According to [13, Section 6], for $i\geq 0$ we say that $h_{i}$ fixes (or reaches fixation) if there exists $t_{0}>0$ such that for all $t>t_{0}$, the trajectory $h_{i}$ is an ancestor (i.e., parent, or parent of parent etc.) of all trajectories $h_{j}$ such that $h_{j}(t)>0$. If the $i$-th trajectory is alone in the system with its parent, $h_{i}$ reaches height 1 at time $t_{i}+\frac{1}{a_{i}}$ and immediately fixes, but due to clonal interference with other trajectories, this is not always the case: The slope of the $i$-th trajectory may change due to resident changes after time $t_{i}$.
Given our algorithm in the proof of Theorem 11 for the resident fitness, for finite $n$ we can now determine in linear time which trajectories fix. Indeed, the trajectories that fix are precisely the ultimate resident and its ancestors, and this ancestral line can be found in $O(n)$ time using linear search. We can search for the parent of the ultimate resident starting from the end of the list in the output of the algorithm, and then we can search for the parent of this parent continuing backwards etc.
5.2 Computing height functions and values of the resident fitness
Given the form of the output of the algorithm in Section 4.2, which we can convert from a doubly linked list into an array at the end of algorithm within $O(n)$ time, for any fixed time $t>0$ we can determine the resident fitness and resident type at time $t$ within a runtime of $O(\log n)$ using binary search. Next, if we want to compute $h_{i}(t)$ for some $i$ (where $t\geq t_{i}$), we can first compute the resident fitness values $f(t_{i})$ and $f(t)$ and the resident types $\varrho(t_{i})$ and $\varrho(t)$ at times $t_{i}$ and $t$. Then if say $s_{j}<t_{i}<s_{j+1}$ and $s_{k}<t<s_{k+1}$, we know that in the continued lines representation, the $i$-th ray $s\mapsto l_{i}(s)$ starts with slope $f(t_{i})+a_{i}$ at height $y_{j}+(t_{i}-s_{j})b_{m_{j}}-1$ at time $t_{i}$. Now, in the original system of interacting trajectories, we have $h_{i}(t)=l_{i}(t)-(l_{\varrho(t)}-1)=l_{i}(t)-(l_{m_{k}}(t)-1)$ if this quantity is positive and $h_{i}(t)=0$ otherwise. Thus, $h_{i}(t)$ can also be computed in $O(\log n)$ time (for fixed $t$). Moreover, the entire piecewise linear trajectory $t\mapsto h_{i}(t)$ can be determined in $O(n)$ time as follows: at time $t_{i}$ it starts with slope $a_{i}$, and afterwards at each resident change time $s_{j}$, its right slope $h_{i}^{\prime}(\cdot+)$ gets reduced by the slope increment of the resident at this time. This way, $h_{i}(s_{j})$ can be computed inductively for all $s_{j}>t_{i}$; when we first obtain a negative number, we check when precisely between $s_{j-1}$ and $s_{j}$ the trajectory $h_{i}$ hits zero, and from that time on we let $h_{i}$ be identically zero.
Acknowledgements
The authors thank Felix Hermann for interesting discussions and comments. He provided the first construction proving Proposition 9, which was somewhat different from the one appearing in our proof.