### v2

# Resident fitness computation in linear time and

Abstract: The notion of a system of interacting trajectories was recently introduced in [16]. Such a system of $[0,1]$-valued piecewise linear trajectories arises as a scaling limit of the system of logarithmic subpopulation sizes in a population-genetic model (more precisely, a Moran model) with mutation and selection. By definition, the resident fitness is initially 0 and afterwards it increases by the ultimate slope of each trajectory that reaches height 1.
We show that although the interaction of $n$ trajectories may yield $\Omega(n^{2})$ slope changes in total, the resident fitness function can be computed algorithmically in $O(n)$ time. Our algorithm uses the so-called continued lines representation of the system of interacting trajectories. In the special case of Poissonian interacting trajectories where the birth times of the trajectories form a Poisson process and the initial slopes are random and i.i.d., we provide a linear bound on the expected total number of slope changes.
Keywords: (Poissonian) interacting trajectories, continued lines representation, algorithmic construction, resident fitness, speed of adaptation.
1 Introduction
The concept of a system of interacting trajectories has been recently introduced in [16]. Such a system consists of continuous and piecewise linear trajectories in a 2D coordinate system, where the $x$ axis represents time and the $y$ axis represents a magnitude that corresponds to the size of subpopulations of a large population of constant total size on a logarithmic scale. The value of the trajectories is in $[0,1]$ at each time. Each trajectory has a birth time and an initial slope. Until the birth time, the trajectory is constant zero, then it starts growing linearly with its initial slope. When a trajectory reaches height 1, it becomes constant 1, while the slope of all other trajectories currently at a positive height is reduced by the ultimate slope of this trajectory before reaching height 1. The slope of each trajectory is nonincreasing, and trajectories that reach height 0 again stay constant 0 forever. This interactive dynamics models the exponential growth/decay of mutant subpopulations in a certain stochastic population-genetic model, more precisely a Moran model, with mutation and selection, under a logarithmic scaling. Here, selection is strong and the mutation rate is in the so-called Gerrish–Lenski regime (dating back to [15]), see Remark 2.4 below.
A special case is the system of Poissonian interacting trajectories (PIT), where birth times are random, forming a Poisson process, and the positive initial slopes are random and i.i.d. (i.e. independent and identically distributed) and independent of the birth times. Under a suitable scaling of parameters, the family of logarithmic subpopulation sizes indeed converges to the PIT in distribution (in a suitable space with a suitable topology, see [16, Theorem 2.7]). In the PIT, the resident fitness $F(t)$ is defined as the sum of the ultimate slope of trajectories reaching height 1 up to time $t$. If the distribution of initial slopes has a finite first moment, $F(t)/t$ converges almost surely to a deterministic number $\overline{v}$ called the speed of adaptation, and it tends to infinity otherwise (cf. [16, Theorem 2.8]). Moreover, if the initial slopes have a finite second moment, then the resident fitness process satisfies a functional central limit theorem (see [16, Theorem 2.12]). The proofs of the latter two assertions are based on a renewal argument.
A closed formula for $\overline{v}$ is only known in the case when the initial slopes are deterministic and constant. In that case, the speed was already computed in [1]; this result was put into the context of the PIT by [16, Proposition 2.10]. On the other hand, one can clearly see the computation of the value of $F(t)/t$ for a large $t$ as a Monte Carlo simulation for the value of $\overline{v}$. In the case when the fitness advantages have finite variance, we also know from the functional central limit theorem that typical fluctuations of $F(t)/t$ around $\overline{v}$ are of order $1/\sqrt{t}$. Thus, in order to approximate $\overline{v}$, it is useful to find an efficient algorithm for the simulation of the resident fitness of (Poissonian) interacting trajectories.
The main result of the present paper, Theorem 3.3 states that there is a deterministic algorithm that computes the resident fitness for all times in the case of a system of $n$ interacting trajectories (with deterministic birth times and initial slopes, where the birth times have been sorted in advance) in $O(n)$ time. The existence of such a linear-time algorithm is a priori not obvious because as we will show in Proposition 3.1 below, the interactions between the trajectories may yield $\Omega(n^{2})$ slope changes in total. Our general conclusion is that computing the resident fitness at all times is in general easier than determining the slopes of all piecewise linear trajectories in all pieces. Our algorithm is described in terms of the so-called continued lines representation, a process introduced in [16] that is in one-to-one correspondence with the PIT but consists of rays (half-lines) instead of piecewise linear broken lines. As Corollary 3.4 asserts, in case all initial slopes are in $\{1,2,\ldots,k\}$ for some $k\in\mathbb{N}$, the piecewise constant slope functions of all trajectories can be computed in $O(kn)$ time based on our main algorithm.
Moreover, our Proposition 3.5 below states that in the case of the Poissonian interacting trajectories, the expected number of slope changes up to time $t$ is $O(t)$ as $t\to\infty$. The proof of the latter assertion is based on a modified version of the aforementioned renewal argument of [16] combined with classical large-deviation estimates. This proof clearly shows that although the worst-case interaction between $n$ trajectories may provide $\Omega(n^{2})$ slope changes in total, scenarios where there are more than a sufficiently large constant times $t$ slope changes up to time $t$ are exponentially unlikely in the Poissonian case. Based on this proof and additional arguments from [16], we derive Corollary 3.7, a strong law of large numbers for the number of slope changes up to time $t$ in the large-$t$ limit.
Let us note that apart from the PIT, there are many other recent examples of piecewise linear limiting processes arising in stochastic population-genetic [9] and population-dynamic [2, 5, 6, 7, 8, 12, 13, 18] models under a logarithmic scaling. These processes originate from models with higher mutation rates than the one of [16], under a different scaling of time. The limiting processes are typically deterministic and their dynamics is also a bit different from the one of the interacting trajectories, thanks to multiple mutations between two given types of individuals and also due to other biological phenomena depending on the model (e.g. horizontal gene transfer, asymmetric competition, dormancy etc.). It is possible that variants or extensions of our algorithm can be used for simulating these limiting processes efficiently, but due to length constraints, we defer such questions to future work.
The rest of this paper is organized in follows. In Section 2 we recall the definition of a system of (Poissonian) interacting trajectories from [16] and we provide a short summary of its biological motivation. In Section 3 we present our results and in Section 4 we prove them. Finally, in Section 5 we briefly discuss some consequences of our assertions and proofs. More concretely, in Section 5.1 we explain some applications of our algorithm in determining ancestral relations and fixation events of trajectories. Using this, in Section 5.2, we recall variants of the Gerrish–Lenski heuristics [1, 15] and contrast their applicability with the one of our linear-time algorithm for multiple purposes. In Section 5.3 we discuss how to determine height functions of the trajectories and fixed-time values of the resident fitness efficiently using our algorithm and data structure. Finally, in Section 5.5 we sketch potential applications of our algorithm in a parameter regime where biological selection is weaker than in our original model, which is called the case of moderate selection.
2 Setting
The definition of a system of (Poissonian) interacting trajectories that we will use in the present paper is a somewhat simplified version of the definition appearing in [16, Sections 2.1 and 3.1] (see Remark 2.7 below for the differences between the two definitions). It is given as follows.
Writing $\mathbb{N}=\{1,2,\ldots\}$ and $\mathbb{N}_{0}=\{0,1,2,\ldots,\}$, for $n\in\mathbb{N}_{0}$ we define $n]=\{1,2,\ldots,n\}$, moreover we put $\infty$$\mathbb{N}$. We denote the space of continuous and piecewise linear trajectories $h$ from $0,\infty)$ to $[0,1]$ such that $h(t)\in[0,1]$ for all $t\geq 0$ by $\mathcal{C}_{\rm PL}$. Each $h\in\mathcal{C}_{\rm PL}$ has at time $t$ a height $h(t)$ and a (right) slope
$v_{h}(t):=\lim\limits_{\delta\downarrow 0}\tfrac{1}{\delta}(h(t+\delta)-h(t)).$ (1)
Assume that for some $n\in\mathbb{N}_{0}\cup\{\infty\}$ we are given a configuration of pairs $(t_{i},a_{i})$ consisting of a birth time $t_{i}$ and an initial slope $a_{i}$
$\beth=((t_{i},a_{i}))_{1\leq i<n+1}\in([0,\infty)\times(0,\infty))^{[n]},$ (2)
[…]
Definition 2.1 (System of interacting trajectories, [16]).
[…]
$\mathbb{H}=(h_{i})_{0\leq i<n+1}\in\big{(}\mathcal{C}_{\rm PL}\big{)}^{[n]\cup\{0\}}$ (3)
result from the following deterministic interactive dynamics on $(\mathcal{C}_{\rm PL})^{[n]\cup\{0\}}$, where we write $v_{i}:=v_{h_{i}}$ for $0\leq i<n+1$ in order to simplify notation:
  • One initial trajectory $h_{0}\in\mathcal{C}_{\rm PL}$ starts at height $h_{0}(0)=1$ with slope $v_{0}(0)=0$. All other trajectories $h_{i}$, $1\leq i<n+1$ start at height $h_{i}(0)=0$ with slope $v_{i}(0)=0$.
  • Trajectories continue with constant slope until the next birth time is reached or one of the trajectories reaches either 1 from below or 0 from above.
  • For $i\geq 1$, at the birth time $t_{i}$ the slope $v_{i}$ of trajectory $h_{i}$ jumps from 0 to $a_{i}$.
  • Whenever at some time $t$ at least one trajectory reaches height $1$ from below, the slopes of all trajectories whose height is in $(0,1]$ at time $t$ are simultaneously reduced by
    $v^{\ast}:=\max\limits\{v_{i}(t-)\mid 0\leq i<n+1\mbox{ such that }h_{i}(t)=1\},$
    i.e. for all $0\leq i<n+1$ with $h_{i}(t)>0$
    $v_{i}(t):=v_{i}(t-)-v^{\ast}.$ (4)
  • Whenever a trajectory at some time $t$ reaches height 0 from above, its slope is instantly set to 0, and this trajectory then stays at height 0 forever.
    Let $e_{i}$ denote the time when this happens to the $i$-th trajectory during the sequential application of the heuristics (where we set $e_{i}=\infty$ if it never happens) where $0\leq i<n+1$. For $t\geq 0$ we say that the $i$-th trajectory is alive at time $t$ if $t_{i}\leq t<e_{i}$.
[…]
Definition 2.2 (Resident change times, resident type, and resident fitness, [16]).
Let $\mathbb{H}$ be as in (3), following the dynamics specified in Definition 2.1.
  • The times at which one of the trajectories $h_{i}$, $i>0$, reaches height $1$ from below will be called the resident change times.
  • For $t>0$ we call
    $\varrho(t)=\varrho^{(\beth)}(t):=\operatorname*{arg\,max}_{i}\{v_{i}(t-)\mid 0\leq i<n+1,h_{i}(t)=1\},$ (5)
    the resident type at time $t$, and we put $\varrho(0)=\varrho^{(\beth)}(0)=0$.
  • With the definition $f(0)=0$, we define the resident fitness $f(t)$, $t\geq 0$, by decreeing that $f$ at any resident change time $r$ has an upward jump with
    $f(r)-f(r-)=\max\{v_{i}(r-)\mid 0\leq i<n+1,\,h_{i}(r)=1\}$ (6)
and remains constant between any two subsequent resident change times.
The following remark is an adaptation of [16, Equations (3.5) and (3.6)] to our situation.
Remark 2.3 ([16]).
  1. 1.
    The resident fitness $f$ obeys
    $f(t)=a_{\varrho(t)}+f(t_{\varrho(t)}),\quad t\geq 0$ (7)
    with $a_{0}:=0$.
  2. 2.
    If $t\geq 0$ is not a resident change time, then $\varrho(t)$ is the unique $i>0$ for which $h_{i}(t)=1$. If $t$ is a resident change time and $h_{i}$ is the only trajectory that reaches height 1 from below at time $t$, then $\varrho(t)=i$. More generally, $\varrho(t)$ in (5) is the fittest of all types that are at height $1$ at time $t$.
  3. 3.
    The resident type $\varrho(t)$ defined in (5) satisfies (with $t_{0}:=0$)
    $\varrho(t)=\max\{i\colon 0\leq i<n+1,\,h_{i}(t)=1\}$ (8)
    (see [16, Remark 3.4, third arXiv version] for an explanation; this remark is not present in the current, fifth version of [16] on arXiv.).
Remark 2.4.
The biological interpretation of this interactive dynamics is as follows (see [16] for more details). We consider a Moran model with mutation and selection. This population-genetic model is a continuous-time Markov chain. The total population size $N$ of the model is fixed, and the interacting trajectories arise as a scaling limit as $N\to\infty$. The height $h_{i}(t)$ of a given trajectory at a given time $t$ corresponds to the logarithmic size of a subpopulation of genetically identical individuals divided by $\log N$ in the Moran model. Thus, the height of a non-extinct subpopulation is always between 0 and 1, and upon extinction, we cut the trajectory at height 0 by definition, so that it cannot become negative.
At time 0, all $N$ individuals belong to the initial resident population and have the same fitness value 0. Beneficial mutations occur in a randomly chosen individual at time $t_{i}\log N$, with a positive fitness advantage (selective advantage) $a_{i}$, for all $1\leq i<n+1$. This selective advantage allows mutant subpopulations to reproduce faster; its strength does not depend on $N$ (this is why selection is called strong). Mutant subpopulations tend to either go extinct rapidly or grow approximately exponentially, hence their logarithmic population size divided by $\log N$ grows approximately linearly and takes $\Theta(\log N)$ time to reach the vicinity of 1.
The per generation mutation rate is $\mu_{N}=\Theta(1/\log N)$, in other words, the model belongs to the so-called Gerrish–Lenski mutation regime. A characteristic property of this regime is that the waiting time between two consecutive mutations is of the same order of magnitude (in our case it is $\Theta(\log N)$) as the time until a mutant subpopulation reaches height 1. Hence, typically multiple mutant families are fighting with each other for survival, which is a phenomenon referred to as clonal interference.
Speeding up time by a factor of $\log N$, we obtain approximately the system $\mathbb{H}(\beth)$; the corresponding convergence result is [16, Theorem 2.7]. Each mutation happens in an individual of the current resident population with high probability. Hence, the limiting trajectory $h_{i}$ starts with slope $a_{i}$ at time $t_{i}$ and keeps this slope until some subpopulation becomes resident (i.e. its trajectory reaches height 1). Upon this resident change, the slope of each subpopulation currently at a positive height drops by the increment of the resident fitness, i.e. the ultimate slope of the trajectory of the new resident before reaching height 1 (cf. (4)). Thus, slopes may become negative, corresponding to an approximately exponential decay of subpopulations that are already unfit compared to the current resident. The time when a trajectory hits zero marks the extinction time of the corresponding mutant subpopulation in the Moran model.
[…]
  • the role of $(t_{i})_{i\in\mathbb{N}}$ is played by the sequence of arrival times $(T_{i})_{i\in\mathbb{N}}$ of a Poisson process $(T_{i})_{i\in\mathbb{N}}$ with intensity $\lambda>0$ (i.e. putting $T_{0}=0$, $T_{i}-T_{i-1}$, $i=1,2,\ldots$ are i.i.d. exponentially distributed random variables with parameter $\lambda$),
  • whereas the role of $(a_{i})_{i\in\mathbb{N}}$ is played by a sequence $(A_{i})_{i\in\mathbb{N}}$ of i.i.d. strictly positive random variables that is independent of the Poisson process $(T_{i})_{i\in\mathbb{N}}$, where the common marginal distribution of each $A_{i}$ on $(0,\infty)$ is denoted by $\gamma$ (i.e. $\gamma$ is a Borel probability measure on $(0,\infty)$ such that for any $x\in\mathbb{R}$, $\gamma((-\infty,x])=\mathbb{P}(A_{1}\leq x)$).
In this case, we call $\mathbb{H}(\beth)$ the system of Poissonian interacting trajectories (or PIT for short) with parameters $(\lambda,\gamma)$, or briefly as the $\mathrm{PIT}(\lambda,\gamma)$. The corresponding trajectories and their (right) slopes will also be denoted by capital letters $H_{i}$ resp. $V_{i}$ (note that they are deterministic functions of the random input $((T_{j},A_{j}))_{j\in\mathbb{N}}$). The resident fitness $F$ that we mentioned in the introduction corresponds to $f$ from Definition 5 in the case of the PIT. The resident type at time $t\geq 0$ in this context will still be denoted by $\varrho(t)$. We will generally exclude conditions that are degenerate according to the following definition.
Definition 2.5.
[…]
  • There exists $t>0$ and three pairwise distinct indices $i,j,k\in[n]\cap\{0\}$ such that $h_{i}(t)=h_{j}(t)=h_{k}(t)=1$.
  • There exists $j\in[n]\cap\{0\}$ such that the extinction time of $h_{j}$, i.e. $\inf\{t>t_{0}\mid h_{j}(t)=0\}$, is a resident change time or equals $t_{i}$ for some $i\in[n]$.
  • There exists $i\in[n]$ such that $t_{i}$ is a resident change time.
Elementary properties of the Poisson process imply the following claim.
Claim 2.6.
Let $\beth=((T_{i},A_{i}))_{i\in\mathbb{N}}$ as above. Then we have that $\mathbb{P}(\beth\text{ is nondegenerate})=1.$
Remark 2.7.
Apart from considering only finitely many trajectories instead of countably many, the main changes in the above definitions compared to their analogues in [16] are the following:
  • We ignore general initial conditions with multiple trajectories at positive heights at time 0, which [16] used in proofs. For simplicity we allow $t_{1}=0$ (and not only $t_{1}>0$).
  • The original definition of the PIT includes trajectories whose slope and height stays zero forever. These describe mutant families going extinct rapidly by chance shortly after the mutation. This possibility is biologically relevant but algorithmically uninteresting.
Remark 2.8 (The continued lines representation).
An alternative, equivalent representation of the PIT, called the continued lines representation, was provided in [16, first arXiv version, Section 5.3]. (The latter section is only included in the first and second version of [16] on arXiv.) This representation can be simulated without any calculations involving already decaying height functions, which turns out to be useful for algorithmic investigations. We describe this representation for a general system of interacting trajectories $\mathbb{H}(\beth)$ as follows.
[…]
Let us assume that $\mathbb{H}(\beth)$ is nondegenerate. Then, for $k\geq 0$, if $j\in\{0,1,\ldots,k\}$ satisfies $l^{r}(t_{k+1})=l_{j}(t_{k+1})$, we say that $j$ is the parent of $k+1$ (note that there is a unique such $j$ that is well-defined under the assumption of nondegeneracy).
3 Results
For a sequence $(\beth^{(n)})_{n\in\mathbb{N}}$ of nondegenerate initial data $\beth^{(n)}=((t_{i}^{(n)},a_{i}^{(n)}))_{1\leq i<n+1}$, writing $\mathbb{H}(\beth^{(n)})=(h_{i}^{(n)})_{0\leq i<n+1}$, we denote by
$k_{n}(\beth^{(n)})$ $=\sum_{t\geq 0\colon t\text{ resident change time of }\mathbb{H}(\beth^{(n)})}\sum_{i=0}^{n}\mathds{1}_{\{v_{i}^{(n)}(t-)\neq v_{i}^{(n)}(t+)\}\cap\{h_{i}^{(n)}(t)>0\}}$
[…]
the total number of kinks (slope changes) in the system of interacting trajectories $\mathbb{H}(\beth^{(n)})$. Then we have the following result.
Proposition 3.1.
There exists a sequence $(\beth^{(n)})_{n\in\mathbb{N}}$ as above such that $k_{n}(\beth^{(n)})=\Omega(n^{2})$.
The proof of Proposition 3.1 can be found in Section 4.1.
Remark 3.2.
Apart from the slope changes corresponding to the definition of $k_{n}$, at each time $t_{i}$, $i=1,\ldots,n$, there is an additional slope change where the slope of the $i$-th trajectory switches from 0 to $a_{i}$. Moreover, at most $n$ trajectories eventually hit 0 again and then their slope drops to 0. Altogether, these are at most $2n$ additional slope changes, and precisely $2n$ unless there are multiple indices $j$ with the same value of $f(t_{j})+a_{j}$ as the ultimate resident. (In the case of the PIT, the latter is only possible if $\mathbb{P}(A_{1}=c)>0$ holds for some $c>0$.) Hence, given Proposition 3.1, ignoring these $O(n)$ additional slope changes will not yield substantial changes in the worst-case estimates of the number of kinks.
[…]
Theorem 3.3.
The algorithmic complexity of finding $t\mapsto\varrho^{(\beth)}(t)$, $t\geq 0$, for any finite, nondegenerate, $red \beth o$f size $n$ is $\Theta(n)$.
The proof of this theorem will be provided in Section 4.2. The proof of the upper bound is constructive: we present an algorithm that computes the array describing the function in $O(n)$ time. The key idea is to use the continued lines representation of the system, which makes it possible to provide such an algorithm using a simple data structure, despite the possibly large number of kinks. As a byproduct, the algorithm also computes the trajectories in the continued lines representation of the system. From these outputs, the resident fitness function $t\mapsto f(t)$ is also easily computable, since it corresponds to the slopes of the trajectories of the residents.
[…]
In the special case when we additionally know that each $a_{i}$ is contained in the set $[k]$ for some $k\geq 1$, the algorithm appearing in the proof of Theorem 3.3 can easily be modified so that it also computes the piecewise slopes of all interacting trajectories (not only the resident fitness or the slopes of the rays of the continued lines representation) in $O(kn)$ time. This fact is reflected by the following corollary.
Corollary 3.4.
Let $k\in\mathbb{N}$. For any finite, nondegenerate, $red \beth o$f size $n$ such that each initial slope $a_{i}$ lies in $[k]$, the family of piecewise constant slopes $(t\mapsto v_{i}(t))_{i\in[n]}$ can be computed via a deterministic algorithm of runtime $O(kn)$.
The proof of Corollary 3.4 can be found in Section 4.3.
Our next proposition shows that if one considers the PIT where the birth times of trajectories form a Poisson process, then even the expected number of kinks will only grow linearly in time. Due to the time-continuous nature of the Poisson process, here one has to work with a notion of kinks (3) slightly different from the one in (3) above; see Remark 3.6 below for a related discussion.
Let $\lambda,\gamma$ as above and let us consider the $\mathrm{PIT}(\lambda,\gamma)$. For $t\geq 0$ we write $N_{t}=|\{i\geq 1\colon T_{i}\leq t\}|$ where $T_{i}$ is the birth time of the $i$-th trajectory (with the aforementioned convention that $T_{0}=0$), then $N_{t}$ is $\mathrm{Poisson}(\lambda t)$ distributed for any $t>0$, and for any $n\in\mathbb{N}$ and $0=t_{0}<t_{1}<\ldots<t_{n}$, $N_{t_{i}}-N_{t_{i-1}}$, $i=1,\ldots,n$ are independent. Define the total number of kinks up to $t$ as
$K_{t}=\sum_{0<s\leq t\colon s\text{ resident change time}}\ \sum_{i\in\mathbb{N}_{0}}\mathds{1}_{\{H_{i}^{\prime}(t-)\neq H_{i}^{\prime}(t+)\}\cap\{H_{i}(t)>0\}}.$
Proposition 3.5.
We have $\mathbb{E}(K_{t})=O(t)$.
The proof of Proposition 3.5 will be carried out in Section 4.4. This proof is based on an extension of the aforementioned renewal argument of [16], combined with classical large-deviation inequalities. The main message of this proof is that the clustering of trajectories that leads to many kinks (like in the construction in the proof of Proposition 3.1) is exponentially unlikely in the number of participating trajectories if these trajectories follow each other directly in the PIT according to their birth times.
Remark 3.6 (Relation between Propositions 3.1 and 3.5).
$K_{t}$ is the number of resident changes up to time $t$ counted with multiplicity: Each resident change is counted as many times as many trajectories change their slopes due to this resident change. Note that putting $\beth^{(n)}=((T_{i},A_{i}))_{i\in[n]}$, the total number of kinks due to the interactions of the first $N_{t}$ trajectories is in general an upper bound for $K_{t}$, i.e.
$K_{t}\leq k_{N_{t}}(\beth^{(N_{t})}),$
the difference $k_{N_{t}}(\beth^{(N_{t})})-K_{t}$ being the number of those kinks in $\mathbb{H}(\beth^{(N_{t})})$ that happen after time $t$. However, it will be apparent from the proof of Proposition 3.5 that following time 0, after a stopping time with finite expectation, a state will be reached where $H_{i}(t)=0$ for all but one $i$. Hence, by the strong Markov property, this state will be visited infinitely often, with i.i.d. waiting times of finite expectation inbetween. Since a trajectory can only have kinks before its height ultimately reaches 0, it follows that the difference between $K_{t}$ and $k_{N_{t}}(\beth^{(N_{t})})$ is of finite order. This way, Proposition 3.5 can indeed be seen as complementary to Proposition 3.1 in the Poissonian case.
Based on the proof of Proposition 3.5 and the one of [16, Theorem 2.8], in Section 4.4 we will also derive the following corollary, which can be seen as a strong law of large numbers for the total number of kinks.
Corollary 3.7.
Let $E_{0}=0<E_{1}<E_{2}<\ldots$ be the renewal extinction times, defined via
[…]
for $i\geq 1$. We have
$\lim_{t\to\infty}\frac{K_{t}}{t}=\frac{\mathbb{E}(K_{E_{1}})}{\mathbb{E}(E_{1})},$
almost surely. The limit on the right-hand side is in $(0,\infty)$ for any choice of $\gamma$.
Remark 3.8.
Corollary 3.7 is formally analogous to [16, Theorem 2.8] on the speed of adaptation: There, under the additional assumption that $\mathbb{E}(A_{1})<\infty$, using the solitary resident change times $L_{0}=0<L_{1}<L_{2}<\ldots$ defined as
[…]
it was shown that the resident fitness $t\mapsto F(t)$ satisfies
$\lim_{t\to\infty}\frac{F(t)}{t}=\frac{\mathbb{E}(F(L_{1}))}{\mathbb{E}(L_{1})}\in(0,\infty),$
almost surely. We will explain this analogy in more detail and exploit it in the proof of Corollary 3.7 below. Note however that a substantial difference between the two results is that if $\mathbb{E}(A_{1})=\infty$, then $\lim_{t\to\infty}F(t)/t=\infty$ a.s., while in Corollary 3.7 the limit is always finite.
Remark 3.9.
When $A_{1}$ is constant and equal to $c>0$, the limit in (3.7) is $2\lambda$. To see this, note that if all $A_{i}$’s are equal to $c$, then any trajectory born after time 0 will start with slope $c$ at height 0, reach slope 0 at the first resident change following its birth and reach slope c-c at the next resident change. This slope will be preserved until the trajectory reaches height 0 (see also Lemma 4.1). Now, by time $E_{1}$, all trajectories born before $E_{1}$ but the one that is resident at time $E_{1}$ must have reached height 0 again. This accounts for 2 slope changes due to resident changes per trajectory, apart from the 0-th trajectory, which started from height 1 with slope 0, and the trajectory that is resident at time $E_{1}$, which ends at height 1 with slope 0; the latter two trajectories have only had one slope change by time $E_{1}$ each. Thus, almost surely, $\mathbb{E}(K_{E_{1}}|N_{E_{1}})=2N_{E_{1}}$, and hence by the tower property, $\mathbb{E}(K_{E_{1}})=2\mathbb{E}(N_{E_{1}}|N_{E_{1}})=2\mathbb{E}(N_{E_{1}})=2\mathbb{E}(\mathbb{E}(N_{E_{1}}|E_{1}))=2\lambda\mathbb{E}(E_{1})$, which implies that the limit in  (3.7) is indeed $2\lambda$. For general distributions $\gamma$, we expect that computing the limit in (3.8) is harder.
As we already anticipated in Remark 2.7[16] studied a variant of the PIT where trajectories have a chance to stay constant equal to zero, which corresponds to rapid extinction of the given mutant subpopulation in the Moran model. There, in the constant-$c$ case, trajectories arrive at rate $\lambda$ but independently with probability $\frac{1}{1+c}$ they stay constant 0 and only with probability $\widetilde{\lambda}:=\frac{\lambda c}{1+c}$ they follow the dynamics described in Section 1. In that setting, $\lim_{t\to\infty}K_{t}/t$ would be $2\widetilde{\lambda}=\frac{2\lambda c}{1+c}$. We note that this is also the only case where an explicit formula for the speed of adaptation is known. We have (in the latter variant of the PIT) that $\lim_{t\to\infty}F(t)/t=\frac{\lambda c^{2}}{1+c+\lambda}$. This result is [16, Theorem 2.10], which in turn originates from [1, Section 3.1]; see Section 5.2 for more details.
4 Proofs
4.1 Lower bound construction for the number of kinks: proof of Proposition 3.1
Proof.
The sequence $(\mathbb{H}(\beth^{(n)})_{n\in\mathbb{N}}$ corresponding to the statement of the proposition is given as follows. For simplicity, we use the continued lines representation introduced in Remark 2.8 (which is in bijection with the representation as a system of interacting trajectories).
For $n\in\mathbb{N}$ and $k\in\{0,1,\ldots,n\}$, we define the $k$-th ray $l_{k}$ corresponding to the continued lines representation of $\mathbb{H}(\beth^{(n)})$ via the equations $l_{0}(t)=1$, $t\geq 0$, and for $k\geq 1$,
$l_{k}(t)=k\Big{(}t-(1-\frac{1}{k}+\frac{k-1}{n^{2}})\Big{)},\qquad t\geq 1-\frac{1}{k}+\frac{k-1}{n^{2}}.$
First we claim that $l_{k}$ and $l_{k+1}$ intersect at time $1+\frac{2k}{n^{2}}$. Indeed, this is clear for $l_{0}(t)=1$ and $l_{1}(t)=t$. Moreover, for $k\geq 1$, $t$ is the time when $l_{k}$ and $l_{k+1}$ intersect if and only if
$k\big{(}t-1+\frac{1}{k}-\frac{k-1}{n^{2}}\big{)}=(k+1)\big{(}t-1+\frac{1}{k+1}-\frac{k}{n^{2}}\big{)},$
which is satisfied if and only if $t=1+\frac{2k}{n^{2}}$. Since $l_{k+1}$ is steeper than $l_{k}$, this implies that
$l_{k+1}(t)>l_{k}(t)\quad\Leftrightarrow\quad t>1+\frac{2k}{n^{2}}.$
We now claim that it is even true that
[…]
Indeed, recall that at a given time $t$, the index of the resident is the index of the ray that is at the highest position among all rays at time $t$ (which implies that the ray has already been started by time $t$, and in the case of a draw, the index with the higher slope is the index of the resident). Now, (13) follows inductively from the fact that 0 is obviously resident at time 0, type $k$ cannot be resident at time $t<1+\frac{2(k-1)}{n^{2}}$ because then $l_{k-1}(t)>l_{k}(t)$, and it cannot be resident at time $t>1+\frac{2k}{n^{2}}$ either because then $l_{k+1}(t)>l_{k}(t)$, but at any time there must be a resident type, so $l_{k}$ must be resident for $\frac{2(k-1)}{n^{2}}<t<\frac{2k}{n^{2}}$ and thus by construction also for $t=\frac{2(k-1)}{n^{2}}$.
Now we claim that the 0-th trajectory of $\mathbb{H}(\beth^{(n)})$ is still alive at time $1+\frac{2(n-1)}{n^{2}}$, equivalently, $l_{0}(1+\frac{2(n-1)}{n^{2}})>l^{r}(1+\frac{2(n-1)}{n^{2}})-1=l_{n}(1+\frac{2(n-1)}{n^{2}})-1$, where we recall from Remark 2.8 that $l^{r}(t)$ is the height of the current resident trajectory at time $t$. Indeed, we have that
$l_{0}(t)=1,\qquad l_{n}(1+\frac{2(n-1)}{n^{2}})=n\big{(}1+\frac{2(n-1)}{n^{2}}-1+\frac{1}{n}-\frac{n-1}{n^{2}}\big{)}=n\big{(}\frac{2n-1}{n^{2}}\big{)}=2-\frac{1}{n}.$
By construction, for $t>1+\frac{2(n-1)}{n^{2}}$, $l_{0}(t)$ is definitely the smallest one among all values $l_{i}(t)$, $i\in\{0,1,\ldots,n\}$. This implies that all other trajectories of $\mathbb{H}(\beth^{(n)})$ are also alive at time $1+\frac{2(n-1)}{n^{2}}$. Since they are all already born (i.e. the corresponding rays of the continued lines representation are already started) before time $1$, it follows that at all resident change times all trajectories of $\mathbb{H}(\beth^{(n)})$ are alive and thus they change their slopes. This yields $k_{n}(\beth^{(n)})=n(n+1)=\Omega(n^{2})$. ∎
4.2 Proof of Theorem 3.3
The algorithm processes births one at a time, in increasing order, while maintaining an ordered list of residents. For trajectory $i$, the parent is found by linear search from the start of the resident list forwards, and whether it becomes resident is determined by linear search from the end of the resident list backwards.
Recall from Remark 2.8 that in the continued lines representation, each trajectory birth is a pair $(t_{i},a_{i})$, where $t_{i}$ is the birth time and $a_{i}$ is the slope relative to its parent.
A trajectory is a tuple $(t_{i},p_{i},b_{i},y_{i})$, where $t_{i}$ is the birth time, $p_{i}$ is the index of the trajectory of the parent, who is the resident at time $t_{i}$, $b_{i}=b_{p_{i}}+a_{i}$ is the slope, and $y_{i}$ is the initial height, which is one less than the height of the parent at time $t_{i}$. This corresponds to a ray starting at point $(t_{i},y_{i})$ with slope $b_{i}$. A residency is a pair $(s_{j},m_{j})$, where $s_{j}$ is the residency start time and $m_{j}$ is the trajectory index. A residency starts when its trajectory intersects the trajectory of the current resident from below, so the residencies form a convex broken line. The $j$-th residency starts at point $(s_{j},y_{m_{j}}+b_{m_{j}}(s_{j}-t_{m_{j}}))$.
Initially, there is one trajectory in the system, $(t_{0}=0,p_{0}=-1,b_{0}=0,y_{0}=1)$, which is the current resident, $(s_{0}=0,m_{0}=0)$. When processing the $i$-th birth, the forward linear search finds the parent of trajectory $i$ in the list of residents, which is the last resident $j$ with $s_{j}\leq t_{i}$. The $i$-th trajectory is $(t_{i},p_{i}=m_{j},b_{i}=(b_{p_{i}}+a_{i}),y_{i}=(y_{p_{i}}+b_{p_{i}}(t_{i}-t_{p_{i}})-1))$. Since the input is ordered by $t_{i}$, the search for the parent can start from the parent of the previous trajectory.
If ray $i$ is below the convex broken line of residents, it does not become a resident. Otherwise, there is an intersection point, which is found by backward linear search. Starting from the last resident in the list, it discards all residency segments from the previous solution below the new ray until the intersection point $(x,y)$ is found and in that case the new residency segment $(x,i)$ is added to the end.
Since both linear searches pass each trajectory at most once, the algorithm has runtime $O(n)$. Note here that as a byproduct, this algorithm also computed the trajectories $(t_{i},p_{i},b_{i},y_{i})$ in the continued lines representation of the system.
The lower bound construction appearing in the proof of Proposition 3.1 (see Section 4.1) shows that for all $n\in\mathbb{N}$ there exists $\beth^{(n)}$ such that $\mathbb{H}(\beth^{(n)})$ exhibits $n$ resident changes. Therefore, the size of the output of any algorithm computing $t\mapsto\varrho^{(n)}(t)$ is $\Omega(n)$ in the worst case, and hence the runtime of any such algorithm is also $\Omega(n)$.
We conclude that the runtime complexity is indeed $\Theta(n)$, which completes the proof of Theorem 3.3.
4.3 Proof of Corollary 3.4
The key step of the proof of Corollary 3.4 is the following lemma.
Lemma 4.1.
[…]
Proof.
Throughout the proof, we will use the following slightly generalized version of [16, Lemma 2.5] (whose proof is analogous to the one of the original lemma). For all $i\neq j$, recalling that
$e_{i}=\inf\{t>t_{i}\colon h_{i}(t)=0\},$
i.e. $e_{i}$ is the first time when the trajectory $h_{i}$ hits 0 again after its birth time $t_{i}$, we have that
$t\mapsto v_{i}(t)-v_{j}(t)\text{ is constant on }[\max\{t_{i},t_{j}\},\min\{e_{i},e_{j}\}).$
Now assume for a contradiction that for some $t\geq 0$ and $i\geq 0$ we have $v_{i}(t)\leq-2k$. Put
$\tau=\inf\{s\geq 0\colon v_{i}(s)\leq-2k\}.$
It is clear that our assumption is equivalent to the assumption that $\tau<\infty$. Recall that $h_{i}$ is continuous and piecewise linear and hence its right derivative $v_{i}$ is right-continuous. Thus, if $\tau<\infty$, then $v_{i}(\tau)\leq-2k$ holds, and thanks to the non-degeneracy assumption, there is a unique $j\in\mathbb{N}_{0}$ such that the $j$-th trajectory becomes resident at time $\tau$.
Using that $v_{j}(s)\leq k$ for all $s$ and that the $j$-th trajectory becomes resident at time $\tau$, it follows that
$\tau\geq t_{j}+\frac{1}{k}.$
Moreover, by construction, under the assumption $\tau<\infty$, $v_{i}(t_{j})-v_{i}(\tau)$ equals the increment of the resident fitness between times $t_{j}$ and $\tau$ (see (4.3) or also [16, Lemma 2.5]), which is precisely equal to $a_{j}\leq k$. It follows that $v_{i}(t_{j})\leq-k$. Thus, if $\tau<\infty$, then $v_{i}(s)\leq-k$ for all $s\in[t_{j},\tau)$. Therefore, using (4.3) we have
$h_{i}(\tau)=h_{i}(t_{j})+\int_{t_{j}}^{\tau}v_{i}(s)\mathrm{d}s<1-k(\tau-t_{j})<1-k\cdot\frac{1}{k}<0.$
This is a contradiction, and hence it follows that $h_{i}$ hits 0 before time $\tau$, i.e. $e_{i}<\tau$ and thus $\tau=\infty$. This finishes the proof of the lemma. ∎
Now we can carry out the proof of Corollary 3.4.
Proof of Corollary 3.4.
Note that if each $a_{i}$ is integer-valued, then so is $v_{i}(t)$ for all $0\leq i<n+1$ and $t\geq 0$. Since at any resident change time, the slope of any trajectory that is currently at a positive height decreases by at least one and, thanks to Lemma 4.1, the slope of any trajectory at any time ranges between (2k1)-(2k-1) and $k$, it follows that the slope of any trajectory can change at most 3k13k-1 times while the trajectory is at a strictly positive height. In Section 5.3 we describe how to compute the height functions $t\mapsto h_{i}(t)$ for all trajectories in $O(kn)$ time, assuming each trajectory undergoes at most $O(k)$ kinks, which completes the proof of this corollary. ∎
Remark 4.2.
[…]
$t_{2}+\frac{1}{k}<t_{1}+\frac{2}{k-1}-\frac{1}{k}+\frac{1}{k}=t_{1}+\frac{2}{k-1},$
until time $t_{2}+\frac{1}{k}$ the 0-th trajectory does not yet reach height zero again. Consequently, at this time, its slope drops by $v_{2}(t_{2}+\frac{1}{k}-)=k$, i.e. it becomes (2k1)-(2k-1).
In the case of the PIT, if $\lambda>0$ and the support of the measure $\gamma$ contains both $k$ and k1k-1 (i.e. $\mathbb{P}(A_{i}=k)$ and $\mathbb{P}(A_{i}=k-1)$ are both positive), the event that the scenario described in the previous paragraph occurs with $(t_{i},a_{i})=(T_{i},A_{i})$ and $h_{i}=H_{i}$, $i=1,2,3$, has positive probability. Using a restart argument based on e.g. the renewal extinction times defined below in Section 4.4, it follows that for such $\gamma$, the $\mathrm{PIT}(\lambda,\gamma)$ exhibits infinitely many trajectories whose slope reaches (2k1)-(2k-1), almost surely. The same is clearly false e.g. if $k$ is even and the support of $\gamma$ is contained in $\{2,4,\ldots,k\}$.
While Lemma 4.1 guarantees that trajectories can never have slopes less than or equal to 2k-2k when all initial slopes $a_{i}$ are in $(0,k]$, slopes in ((2k1),2k)(-(2k-1),-2k) can be obtained for certain choices of the possible values of $a_{i}$ (using an analogue of the construction in the first paragraph of the present remark). Moreover, since the slope decrement due to one kink can be arbitrarily small, having a finite lower bound on the minimal slope does not imply in general that the number of kinks of a given trajectory is bounded by a constant times the absolute value of the minimum possible slope. This is an obstacle to generalizing Corollary 3.4 to this case.
An analogue of Corollary 3.4 holds whenever the possible initial slopes $a_{i}$ are in $\{\alpha,2\alpha,\ldots,k\alpha\}$ for some $\alpha>0$.
4.4 Proof of Proposition 3.5 and Corollary 3.7
Proof of Proposition 3.5.
Recall the renewal extinction times $E_{i}$ from (3.7). Note that for $i\geq 1$, on the time interval $(E_{i-1},E_{i}]$ only trajectories born in $(E_{i-1},E_{i}]$ can become resident, and only the height functions of the same trajectories plus the one of the trajectory that was resident at time $E_{i-1}$ can suffer any kinks on this time interval due to resident changes. Indeed, any other trajectory born on $0,E_{i-1}]$ goes extinct by the time $E_{i-1}$. Hence, we have $K_{E_{i}}-K_{E_{i-1}}\leq(N_{E_{i}}-N_{E_{i-1}})(N_{E_{i}}-N_{E_{i-1}}+1)$, and for $t\geq 0$, putting $\eta_{t}=|\{i\geq 1\colon E_{i}\leq t\}|=\sup\{i\geq 0\colon E_{i}\leq t\}$, we have $E_{\eta_{t}}\leq t<E_{\eta_{t}+1}$. Note also that $\eta_{t}\leq N_{t}$ for all $t$ because at each $E_{i}$ a trajectory born before $E_{i}$ becomes resident (and one trajectory can only become resident at most once).
It can be proven analogously to the proof of [16, Lemma 5.2] that $E_{i}-E_{i-1}$ is stochastically dominated by a geometric random variable, and the random variables $(E_{i}-E_{i-1})_{i\geq 1}$ are i.i.d. (in other words, $(E_{n})_{n\in\mathbb{N}}$ is a renewal process). They have a finite first moment $m_{1}:=\mathbb{E}(E_{1})$ and a finite second moment $m_{2}:=\mathbb{E}(E_{1}^{2})$. This together with the tower property implies that
$\mathbb{E}(K_{E_{i}}-K_{E_{i-1}})$ $\leq\mathbb{E}((N_{E_{i}}-N_{E_{i-1}})(N_{E_{i}}-N_{E_{i-1}}+1))$
$=\mathbb{E}(\mathbb{E}((N_{E_{i}}-N_{E_{i-1}})(N_{E_{i}}-N_{E_{i-1}}+1)|E_{i}-E_{i-1}))$
$=\mathbb{E}(\lambda^{2}(E_{i}-E_{i-1})^{2}+2\lambda(E_{i}-E_{i-1}))=\lambda^{2}m_{2}+2\lambda m_{1}.$
Thus, for $n\geq 1$,
$\mathbb{E}(K_{E_{n}})=\sum_{i=1}^{n}\mathbb{E}(K_{E_{i}}-K_{E_{i-1}})\leq n(\lambda^{2}m_{2}+2\lambda m_{1}).$
The same arguments also imply the (much less tight) bound
$K_{t}\leq N_{t}(N_{t}+1)$
almost surely for all $t\geq 0$. Since $(E_{n})_{n\in\mathbb{N}}$ is a renewal process with increments having finite mean $m_{1}$, we have
$\lim_{n\to\infty}\frac{E_{n}}{n}=m_{1}$
almost surely thanks to the strong law of large numbers, and
$\lim_{t\to\infty}\frac{\eta_{t}}{t}=\frac{1}{m_{1}}$
almost surely (cf. [11, Theorem 2.5.10] for the latter assertion). Using Hölder’s inequality, (4.4) and and (4.4), we conclude that for $\delta>0$, for all $t$ sufficiently large
$\mathbb{E}(K_{t})$ $\leq\mathbb{E}(K_{E_{\eta_{t}+1}}\mathds{1}_{\{\eta_{t}+1\leq(\frac{1}{m_{1}}+\delta)t\}})+\mathbb{E}(K_{t}\mathds{1}_{\{\eta_{t}+1>(\frac{1}{m_{1}}+\delta)t\}})$
$\leq\mathbb{E}(K_{E_{\lceil(\frac{1}{m_{1}}+\delta)t\rceil}})+\sqrt{\mathbb{E}(K_{t}^{2})}\sqrt{\mathbb{E}\Big{(}(\mathds{1}_{\{\eta_{t}+1>(\frac{1}{m_{1}}+\delta)t\}})^{2}\Big{)}}$
$\leq(\frac{1}{m_{1}}+\delta+1)t(\lambda^{2}m_{2}+2\lambda m_{1})+\sqrt{\mathbb{E}(K_{t}^{2})}\sqrt{\mathbb{P}\Big{(}\eta_{t}+1>(\frac{1}{m_{1}}+\delta)t\Big{)}}$
$\leq(\frac{1}{m_{1}}+\delta+1)t(\lambda^{2}m_{2}+2\lambda m_{1})+\sqrt{\underbrace{\mathbb{E}(N_{t}^{4}+2N_{t}^{3}+N_{t}^{2})}_{\in O(t^{4})}}\sqrt{\mathbb{P}\Big{(}\eta_{t}>(\frac{1}{m_{1}}+\delta)t-1\Big{)}}$
$\leq(\frac{1}{m_{1}}+\delta+1)t(\lambda^{2}m_{2}+2\lambda m_{1})+O(t^{2})\sqrt{\mathbb{P}\Big{(}\eta_{t}>(\frac{1}{m_{1}}+\delta)t-1\Big{)}}.$
Here, the assertion that $\mathbb{E}(N_{t}^{4}+2N_{t}^{3}+N_{t}^{2})\in O(t^{4})$ follows from the well-known fact that a Poisson distributed random variable $X$ with parameter $\lambda$ satisfies $\mathbb{E}(X(X-1)\ldots(X-k+1))=\lambda^{k}$ for any $k\geq 1$, from which one can easily derive that $\mathbb{E}(X^{4})=(1+o(1))\lambda^{4}$ as $\lambda\to\infty$.
To treat the last term on the right-hand side of (4.4), we use standard large-deviation estimates. First note that there exists $\delta^{\prime}>0$ not depending on $t$ such that
$\Big{\{}\eta_{t}>\big{(}\frac{1}{m_{1}}+\delta\big{)}t-1\Big{\}}=\Big{\{}E_{\lceil(\frac{1}{m_{1}}+\delta)t\rceil-1}\leq t\Big{\}}\subseteq\Big{\{}E_{\lceil(\frac{1}{m_{1}}+\delta)t\rceil-1}\leq(1-\delta^{\prime})m_{1}\big{(}\lceil(\frac{1}{m_{1}}+\delta)t\rceil-1\big{)}\Big{\}}.$
Since $\Lambda(\alpha):=\log\mathbb{E}(\mathrm{e}^{\alpha E_{1}})<\infty$ for $\alpha>0$ small (which also follows from the stochastic domination by a geometric random variable), Cramér’s theorem [10, Section 2.2] implies that if $\delta^{\prime}$ is sufficiently small, then
$\limsup_{n\to\infty}\frac{1}{n}\log\mathbb{P}\Big{(}E_{n}\leq(1-\delta^{\prime})m_{1}n\Big{)}\leq-\inf_{t\leq(1-\delta^{\prime})m_{1}}\Lambda^{*}(t)=:-I_{0}<0,$
where the rate function $\Lambda^{*}\colon\mathbb{R}\to[0,\infty]$ is given as
$\Lambda^{*}(a)=\sup_{\alpha\in\mathbb{R}}(\alpha a-\Lambda(\alpha)),$
and together with the fact that $\Lambda^{\prime}(0)=m_{1}$ and $\Lambda^{*}(m_{1})=0$, the strict negativity in (4.4) follows from the assertion of [10, Exercise 2.2.24] for $\delta^{\prime}>0$ small enough. We conclude that for all sufficiently large $t$,
$\sqrt{\mathbb{P}\Big{(}\eta_{t}>(\frac{1}{m_{1}}+\delta)t\Big{)}}\leq\mathrm{e}^{-\frac{1}{2}I_{0}(\lceil(\frac{1}{m_{1}}+\delta)t\rceil-1)}\leq\mathrm{e}^{-\frac{1}{2}I_{0}(\frac{1}{m_{1}}+\delta/2)t}.$
Thus, the right-hand side of (4.4) is bounded from above by
$(\frac{1}{m_{1}}+\delta+1)t(\lambda^{2}m_{2}+2\lambda m_{1})+O(t^{2})\mathrm{e}^{-\frac{1}{2}I_{0}(\frac{1}{m_{1}}+\delta/2)t}\in O(t),$
as wanted. ∎
Proof of Corollary 3.7.
The analogy between Corollary 3.7 and [16, Theorem 2.8] that we anticipated in Remark 3.8 is the following. $(E_{i})_{i\geq 1}$ is a renewal process (see the proof of Proposition 3.5 for details), just as $(L_{i})_{i\geq 1}$, and
$\widehat{K}_{t}=\sum_{i\geq 0\colon E_{i}\leq t}K_{E_{i}}-K_{E_{i-1}}=\sum_{i\geq 0}K_{E_{i}}\mathds{1}_{[E_{i},E_{i+1})}(t)$
is a renewal reward process with renewal times $E_{i}$, i.e. it has i.i.d. increments at the renewal times $E_{i}$ that are independent of the $E_{i}$’s, analogously to the process
$\widehat{F}(t)=\sum_{i\geq 0\colon L_{i}\leq t}F(L_{i})-F(L_{i-1})=\sum_{i\geq 0}F(L_{i})\mathds{1}_{[L_{i},L_{i+1})}(t)$
defined in [16, Section 5.2], which is a renewal reward process with renewal times $L_{i}$. For $t\geq 0$ let $m(t)$ be such that $E_{m(t)}\leq t<E_{m(t)+1}$. Then we have
$\frac{K_{E_{m(t)}}/m(t)}{E_{m(t)+1}/m(t)}\leq\frac{K_{t}}{t}\leq\frac{K_{E_{m(t)+1}}/m(t)}{E_{m(t)}/m(t)},$
analogously to [16, Equation (5.9)]. Since
  • $m(t)\to\infty$ as $t\to\infty$,
  • $E_{m}$ is a sum of $m$ i.i.d. copies of $E_{1}$ which have finite (and clearly strictly positive) expectation (cf. the proof of Proposition 3.5),
  • and $K_{E_{m}}$ is a sum of $m$ i.i.d. copies of $K_{E_{1}}$,
both the left- and the right-hand side of (4.4) converge almost surely to $\mathbb{E}(K_{E_{1}})/\mathbb{E}(E_{1})$. Hence, in order to finish the proof of the corollary, it suffices to show that $\mathbb{E}(K_{E_{1}})\in(0,\infty)$. The fact that $\mathbb{E}(K_{E_{1}})<\infty$ follows from (4.4) for $i=1$. Finally, $\mathbb{E}(K_{E_{1}})$ is positive because $K_{E_{1}}$ is almost surely at least 2. Indeed, if the input of the PIT is nondegenerate, then between times 0 and $E_{1}$, the slope of the initial resident trajectory must become negative, while the slope of the resident at time $E_{1}$ must become zero. ∎
5 Discussion
5.1 Ancestral relations and fixations
According to [16, Section 6.1], for $i\geq 0$ we say that $h_{i}$ fixes (or reaches fixation) if there exists $t_{0}>0$ such that for all $t>t_{0}$, the trajectory $h_{i}$ is an ancestor (i.e., parent, or parent of parent etc.) of all trajectories $h_{j}$ such that $h_{j}(t)>0$. If the $i$-th trajectory is alone in the system with its parent, $h_{i}$ reaches height 1 at time $t_{i}+\frac{1}{a_{i}}$ and immediately fixes. But when other trajectories are present, due to clonal interference with them, this is not always the case: The slope of the $i$-th trajectory may change due to resident changes after time $t_{i}$.
Given our algorithm in the proof of Theorem 3.3 for the residents, for finite $n$ we can now determine in linear time which trajectories fix. Indeed, the trajectories that fix are precisely the ultimate resident and its ancestors, and this ancestral line can be found in $O(n)$ time using linear search. We can search for the parent of the ultimate resident starting from the end of the list in the output of the algorithm, and then we can search for the parent of this parent continuing backwards etc.
5.2 Relation to Gerrish–Lenski type heuristics
The Gerrish–Lenski heuristics [15] and its refined version [1] provide a heuristic prediction regarding which mutations reach fixation in the population. In the context of the PIT, using these heuristics one can estimate the speed of adaptation explained in Section 1. Here we provide an informal introduction of these heuristics and their purposes, we discuss some issues with their applicability, and we explain how our linear-time algorithm for the residents can help resolve such issues. For a more formal summary of the heuristics, we refer the reader to [16, Section 6.2].
Recall the notion of solitary resident changes from Remark 3.8 and call a trajectory who becomes resident at a solitary resident change time a solitary resident. This notion easily generalizes to any system of interacting trajectories with nondegenerate input. Consider a nondegenerate $\beth=((t_{i},a_{i}))_{1\leq i<n+1}$ with $n\in\mathbb{N}_{0}\cap\{\infty\}$. In this context, the refined Gerrish–Lenski heuristics [1] asserts that for $1\leq i<n+1$, the $i$-th trajectory of the system of interacting trajectories $\mathbb{H}(\beth)$ born at time $t_{i}$ with initial slope $a_{i}$ becomes solitary resident at time $t_{i}+\frac{1}{a_{i}}$ unless
  1. (i)
    there exists some $1\leq j<i$ such that $t_{j}<t_{i}<t_{j}+\frac{1}{a_{j}}$ and $a_{j}\geq a_{i}$, or
  2. (ii)
    there exists some $i<j<n+1$ such that $t_{i}<t_{j}<t_{i}+\frac{1}{a_{i}}$ and $a_{j}>a_{i}$,
and otherwise it never becomes resident. If situation (i) occurs, then we call the $j$-th trajectory a killer of the $i$-th trajectory from the past, while in situation (ii) we call the $j$-th trajectory a killer of $i$-th trajectory from the future. The original Gerrish–Lenski heuristics [15] asserts that the $i$-th trajectory of $\mathbb{H}(\beth)$ becomes solitary resident at time $t_{i}+\frac{1}{a_{i}}$ unless (ii) occurs and otherwise it never becomes resident. The relation between (solitary) resident changes and fixation (cf. Section 5.1) is that only trajectories that become resident can fix and being an ancestor of a solitary resident trajectory is equivalent to eventual fixation. See [16, Section 6.1] for a justification and further details. Note that if all $a_{i}$’s are the same, then there are no killers from the future and every resident change is solitary.
Applying these heuristics sequentially over the birth times, one can guess which trajectories become (solitary) resident. One first asks if the first trajectory becomes resident, if the heuristics says no, one discards the first trajectory and moves on to the second trajectory etc., taking only non-discarded trajectories into account as potential killers from the past. It was implicitly already included in [1] and later also in [16] that if all $A_{i}$’s are equal, then this sequential application of the refined Gerrish–Lenski heuristics provides the correct set of indices of trajectories becoming (solitary) resident. This procedure yields the correct formula for the speed of adaptation originating from the thinning heuristics introduced in [1, Section 3.1], which was described in terms of the PIT in [16, Proposition 2.10]. On the other hand, in the following example we argue that if the $A_{i}$’s are not almost surely constant, then the sequential application of the refined Gerrish–Lenski makes errors almost surely. Our argument is a slight generalization of the example included in [16, Figure 7], illustrated by Figure 1.
Example 5.1.
Assuming that the $A_{i}$’s are not almost surely constant, for any fixed $\varepsilon>0$ there exist some real numbers $a,b$ such that $0<3\varepsilon<a<b<b+\varepsilon$ and $\mathbb{P}(A_{1}\in(a-\varepsilon,a]),\mathbb{P}(A_{2}\in[b,b+\varepsilon)),\mathbb{P}(A_{3}\in(b-\varepsilon,b])$ are all positive. (E.g. the case when $\gamma$ is a discrete measure supported on two points: $a$ and $b$, we choose $A_{1}=a$ and $A_{3}=A_{2}=b$ for any $\varepsilon$.) Now, with positive probability we have $a-\varepsilon<A_{1}\leq a$, $b-\varepsilon<A_{3}\leq b\leq A_{2}<b+\varepsilon$, $T_{1}<T_{2}<T_{1}+\frac{1}{A_{1}}<T_{3}<T_{2}+\frac{1}{A_{2}}<T_{3}+\frac{1}{A_{3}}<T_{4}$, and
$t^{\prime}:=T_{1}+\frac{1}{A_{1}}+\frac{1-(T_{1}+\frac{1}{A_{1}}-T_{2})A_{2}}{A_{2}-A_{1}}>T_{3}+\frac{1}{A_{3}}.$
Then, since $T_{2}+\frac{1}{A_{2}}>T_{1}+\frac{1}{A_{1}}$, the first trajectory becomes resident at time $T_{1}+\frac{1}{A_{1}}$, so that the slope of the second trajectory gets reduced to $A_{2}-A_{1}\leq b+\varepsilon-a+\varepsilon=b-a+2\varepsilon$. Now, we have $0<b-a+2\varepsilon<b-\varepsilon\leq A_{3}$, and since condition (5.1) holds, at time $T_{3}+\frac{1}{A_{3}}$ the third trajectory becomes resident and thus the second one, which would become resident at time $t^{\prime}$ if there were no further kinks after time $T_{1}+\frac{1}{A_{1}}$, never becomes resident. Here, we also used that $T_{4}>T_{3}+\frac{1}{A_{3}}$ and thus the slope of the third trajectory stays equal to $A_{3}$ until the trajectory becomes resident. Thus, the first and the third trajectory fix eventually, while the second does not. Now, since $A_{2}>A_{1}$ and $T_{1}<T_{2}<T_{1}+\frac{1}{A_{1}}$, the second trajectory is a killer of the first one from the future, and since $A_{2}\geq A_{3}$ and $T_{2}<T_{3}+\frac{1}{A_{3}}$, it is also a killer of the third one from the past. On the other hand, the second trajectory has no killer from its past or future (where we also used that $T_{2}+\frac{1}{A_{2}}<T_{4}$). Thus, a sequential application of the refined Gerrish–Lenski heuristics provides the wrong prediction that the second trajectory becomes resident at time $T_{2}+\frac{1}{A_{2}}$, and that the first and the third one never become resident.
[…]
Figure 1: Illustration of Example 5.1 in the case $A_{2}=A_{3}$. The smooth lines represent the true fate of the trajectories of the PIT. The dashed lines show how the zeroth (orange), second (red) and third (green) trajectory would evolve after time $T_{1}+1/A_{1}$ in absence of the first (blue) one. The sequential application of the heuristics suggests this evolution since the first trajectory has a killer in its future and hence it gets discarded. The slope of the dashed green trajectory becomes zero at time $T_{2}+1/A_{2}$ (since $A_{3}=A_{2}$; in the case $A_{3}<A_{2}$ it would turn negative). Without the additional kink at time $T_{3}+1/A_{3}$, the smooth red trajectory would continue as the dotted one, reaching height 1 at time $t^{\prime}$ introduced in (5.1).
Since whether such a configuration of trajectories occurs between two consecutive solitary resident changes is independent of whether it occurs between other consecutive solitary resident changes, the time until the first occurrence of the event can be bounded by a geometrically distributed random variable and it is therefore almost surely finite. We conclude that almost surely, the sequential application of the refined Gerrish–Lenski heuristics leads to errors infinitely often.
Remark 5.2.
Under similar assumptions on the birth times, in [16, Figure 7] it was assumed that $A_{1}=a,A_{2}=b$ and $A_{3}=c$ satisfy $a,c<b$ and $a+c>b$. This leads to similar conclusions but does not cover the case when $\gamma$ is a discrete measure supported on two points (there we need to choose $c=b$). Further, in Example 5.1 we allowed for $\varepsilon$-errors to cover the case when $\gamma$ has less than two atoms.
Based on Example 5.1, for non-constant initial slopes we recommend using our linear-time algorithm instead of the (refined) Gerrish–Lenski heuristics for determining the resident fitness and the sequence of resident trajectories. It yields a correct result and it increases the runtime by at most a constant factor.
We nevertheless believe that other applications of Gerrish–Lenski type heuristics can still be useful. While our algorithm is suitable for providing a Monte Carlo estimate on the speed of adaptation in the PIT (cf. Section 1) and similarly also on the distribution of the resident change times or of the initial fitnesses of residents or fixing trajectories, finding explicit formulas for these quantities still seems difficult for a general $\gamma$. Unlike the original Gerrish–Lenski heuristics, the refined one already seems to provide quite a good numerical approximation for the speed of adaptation (cf. [16, Figure 6]) given in terms of simple explicit integrals (see [16, Section 6.2]), although it tends to underestimate the speed slightly. Further refinements of the heuristics taking into account not only pairwise but also higher-order interactions between trajectories and distinguishing between solitary residency and fixation may improve the precision and help understand better how clonal interference slows down adaptation.
5.3 Computing the height function
Recall that $t\mapsto\varrho^{(\beth)}(t)$ specifies the resident type at time $t\geq 0$, and its array representation can be computed in a runtime of $O(n)$ as in Section 4.2. This algorithm also produces all trajectories in the continued lines representation, each given by a tuple $(t_{i},p_{i},b_{i},y_{i})$ with birth time $t_{i}$, parent index $p_{i}$, slope $b_{i}$, and initial height $y_{i}$. Using the output of this algorithm, the following computations can be made.
The $i$-th ray in the continued lines representation is $l_{i}(t)=y_{i}+b_{i}(t-t_{i})$ for $t\geq t_{i}$. From this, in the original system of interacting trajectories, we have $h_{i}(t)=l_{i}(t)-(l_{p_{i}}(t)-1)$ if this quantity is positive and $h_{i}(t)=0$ otherwise. This can be evaluated at any given $t\geq 0$ in $O(1)$.
Moreover, the entire piecewise linear trajectory $t\mapsto h_{i}(t)$ can be determined in $O(n)$ time as follows: at time $t_{i}$ it starts with slope $a_{i}$, and afterwards at each resident change time $s_{j}$, its right slope $v_{i}(\cdot)$ is reduced by the slope increment of the resident at this time. This way, $h_{i}(s_{j})$ can be computed iteratively from the residency array for all $s_{j}>t_{i}$; when we first obtain a negative number, we check when precisely between $s_{j-1}$ and $s_{j}$ the trajectory $h_{i}$ hits zero, and from that time on we let $h_{i}$ be identically zero.
Finally, to compute $h_{i}(t)$ for all $i$, one may apply the previous procedure independently for each $i$, resulting in an $O(n^{2})$ algorithm. If it is known in advance that each trajectory undergoes at most $O(k)$ kinks, a more efficient approach is to merge the residency array (keyed by residency start time) with the trajectory array (keyed by birth time) and, at each residency change, keep track of and update only the trajectories currently alive. As each trajectory experiences at most $O(k)$ kinks, this yields an algorithm of runtime $O(kn)$.
[…]
For a lower bound, consider an unordered list of pairwise distinct integers $\{x_{1},\dots,x_{n}\}$. For $\beth=\{(x_{1},2),\dots,(x_{n},2)\}$, all trajectories become resident in increasing order of $x_{i}$. Thus, any algorithm that computes the array representation of the piecewise constant functions $t\mapsto\varrho^{(\beth)}(t)$ or $t\mapsto f(t)$ implicitly sorts the $x_{i}$, which in the comparison model requires $\Omega(n\log n)$ time.
Furthermore, the parent of each trajectory $(x_{i},2)$ is the one immediately preceding it in the sorted order of the $x_{i}$. Selecting the maximum $x_{i}$ and repeatedly taking the parent recovers the entire parent–child chain, yielding the sorted sequence in strictly decreasing order of the $x_{i}$. This means that computing the parent-child relationships between the trajectories is as least as hard as sorting. In the comparison model this establishes a lower bound of $\Omega(n\log n)$.
[…]
5.5 Potential extension of the algorithm to the case of moderate selection
Consider a variant of the Moran model where for some $b\in(0,1)$ fixed, the random fitness advantages $A_{i}$ are multiplied by the vanishing factor $N^{-b}$ depending on the total population size $N$. In this setting, mutants typically survive initial fluctuations with probability of order $N^{-b}$ (cf. [3, 4] for more precise statements in this direction). It was conjectured in [16, Section 7.2] based on simulations and heuristic computations that if one still chooses the per generation mutation rate as $\Theta(1/\log N)$ and speeds up time by a factor of $N^{b}\log N$, the system of rescaled logarithmic frequencies of mutants surviving the initial fluctutations converges to a version of the PIT where trajectories start from height $b$ instead of 0 upon birth and become zero after reaching height $b$ from above again (and otherwise they follow the dynamics described in Section 2). We note that the case when fitness advantages are proportional to $N^{-b}$ for some $b\in(0,1)$ is generally referred to as the case of moderate selection, cf. [3, 4].
In the variant of the continued lines representation corresponding to this variant of the PIT, rays are born 1b1-b (instead of $1$) below the current height of the resident ray, and reaching the current resident height minus (1b1-b) again corresponds to becoming 0 again. It is clear that our algorithm for determining the resident fitness remains valid for this variant of the continued lines representation, still with linear runtime and with the height of newborn trajectories being computed accordingly.
Acknowledgements
The authors thank Felix Hermann for interesting discussions and comments. He provided the first construction proving Proposition 3.1, which was somewhat different from the one appearing in our proof. In the context of Remark 4.2, he was the first to observe that a slope of (2k1)-(2k-1) occurs in the case $k=2$. The authors also thank Anton Wakolbinger for historical remarks and comments on the Gerrish–Lenski type heuristics, which we relied on in the discussion of Section 5.2.